### v1

# Scientific Large Language Models: A Survey on Biological & Chemical Domains

Abstract.
Large Language Models (LLMs) have emerged as a transformative power in enhancing natural language comprehension, representing a significant stride toward artificial general intelligence. The application of LLMs extends beyond conventional linguistic boundaries, encompassing specialized linguistic systems developed within various scientific disciplines. This growing interest has led to the advent of scientific LLMs, a novel subclass specifically engineered for facilitating scientific discovery. As a burgeoning area in the community of AI for Science, scientific LLMs warrant comprehensive exploration. However, a systematic and up-to-date survey introducing them is currently lacking. In this paper, we endeavor to methodically delineate the concept of “scientific language”, whilst providing a thorough review of the latest advancements in scientific LLMs. Given the expansive realm of scientific disciplines, our analysis adopts a focused lens, concentrating on the biological and chemical domains. This includes an in-depth examination of LLMs for textual knowledge, small molecules, macromolecular proteins, genomic sequences, and their combinations, analyzing them in terms of model architectures, capabilities, datasets, and evaluation. Finally, we critically examine the prevailing challenges and point out promising research directions along with the advances of LLMs. By offering a comprehensive overview of technical developments in this field, this survey aspires to be an invaluable resource for researchers navigating the intricate landscape of scientific LLMs.
Scientific domain, large language models, protein, molecule, genome
copyright: acmcopyrightjournalyear: 2024ccs: Computing methodologies Natural language processingccs: Applied computing Computational biology
1. Introduction
“The limits of my language mean the limits of my world.”
Ludwig Wittgenstein
Humanity acquires knowledge of the world through perception and cognition, where natural languages (i.e., human languages) stand as the quintessential medium for articulating this world knowledge. Historically, this plethora of world knowledge has been expressed, chronicled, and disseminated through natural languages. Currently, Large Language Models (LLMs) stand as cutting-edge tools in processing natural language and gathering world knowledge. Typically, LLMs refer to those based on Transformer-based architectures with hundreds of millions (or even billions) of trainable parameters, trained on extensive textual corpus (Shanahan, 2022). Typical examples include GPT-3 (Brown et al., 2020), PaLM (Chowdhery et al., 2023), Galactica (Taylor et al., 2022a), LLaMA (Touvron et al., 2023b), ChatGLM (Zeng et al., 2023a) and Baichuan 2 (Baichuan, 2023). They have exhibited strong capacities to understand natural language and address complex tasks (in the manner of text generation), and have incited substantial interest in both academic and industrial domains. The exceptional performance of LLMs sparks hope that they may evolve into Artificial General Intelligence (AGI) in our current era.
Besides natural languages, to encapsulate more specialized science knowledge, an assortment of scientific languages has been developed, as illustrated in Fig. 1. This encompasses textual expressions in the scientific research domains, mathematical languages to define mathematical formulas, chemical languages such as SMILES that represent molecular structures, and biological languages that describe proteins or genomes, and detail the complex constitution of living organisms. These scientific languages come with their distinct vocabularies, where each term holds a specific meaning that can be completely different from natural languages. For example, the character “C” in English represents the amino acid Cysteine in protein languages (GDR et al., 1984), while in the SMILES language system, it denotes a Carbon atom (Weininger, 1988). Furthermore, experts in specific domains establish grammatical rules to organize these terms, enabling the construction of sentences with precise semantic functions. For instance, computational chemists create grammatical rules ensuring the accuracy of machine-generated molecules in SELFIES format (Krenn et al., 2020). After decades of evolution, scientific languages have become invaluable tools, significantly accelerating scientific discoveries. Due to the potential semantic and grammatical differences between scientific and natural languages, existing general LLMs (such as ChatGPT  [^1] https://chat.openai.com or GPT-4 (OpenAI, 2023)) often fail to properly deal with scientific data like molecules and proteins (AI4Science and Quantum, 2023). As the well-known Austrian philosopher Ludwig Wittgenstein indicates, “The limits of my language mean the limits of my world.” (Ramsey, 1923) The world of general LLMs can be limited to natural languages.
To facilitate the understanding of scientific languages, researchers have devised Scientific Large Language Models (Sci-LLMs) customized for various scientific domains and disciplines. For instance, molecular language models have been developed to represent molecule structures as a string of atoms and chemical bonds (Li and Jiang, 2021). These models aid in predicting molecular properties (Wang et al., 2019), designing new drugs (Zhang et al., 2022b), and proposing retrosynthesis routes (Schwaller et al., 2021). Similarly, protein language models operate based on sequences of amino acids (Rives et al., 2019; Brandes et al., 2022). They are used to forecast 3D protein structures and functions (Lin et al., 2022), enhance existing proteins for improved fitness (Notin et al., 2023b), and create new proteins with specific functionalities (Nijkamp et al., 2023). As a burgeoning area within the AI-for-Science research, many Sci-LLMs have been proposed with modified architectures, learning methods, training corpus, and evaluation benchmarks and criteria. Despite their notable achievements, these models have mostly been explored within their respective research domains. There is currently a lack of a comprehensive review that unifies these language modeling advancements.
In this survey, we set out to fill this gap by systematically reviewing the technical advancement of Sci-LLMs with a close reference to general LLMs. Considering the expansive scope of scientific languages, we focus our investigation on biological and chemical languages. Specifically, our examination encompassed molecular languages, protein languages, and genomic languages. In addition to these specialized scientific languages, we recognize the immense scientific knowledge embedded in textbooks, patents, and research papers, composed in natural languages. Consequently, we explore the textual LLMs with an emphasis on scientific knowledge, and more importantly, the multimodal LLMs that encompass various types of scientific languages.
When delving into each language system, we first review the LLM architectures and categorize them into three classes: encoder-only, decoder-only and encoder-decoder. Then we report the model capabilities and summarize the typical downstream tasks that Sci-LLMs can conduct. In terms of model training and evaluation, we collect a bunch of commonly used training corpus and evaluation benchmarks. Finally, we present the proper criteria for discriminative and generative tasks of scientific language modeling.
Figure 1. Illustrations that general LLMs struggle to effectively handle scientific languages, such as molecules, RNA and amino acid sequences in this example.
This survey is confined within specific boundaries. Firstly, we focus on scientific languages, specifically chemical and biological languages. We have excluded languages lacking both universally defined vocabularies and grammatical structures, such as mathematical languages. Secondly, when discussing textual LLMs, our emphasis remains on chemical and biological domain knowledge expressed in natural languages. This choice ensures a consistent and coherent interaction with languages specific to chemistry and biology, such as molecular and protein languages. Thirdly, our technical exploration is primarily confined to Transformer-based language models. We have not included alternative neural architectures like graph neural networks and diffusion models, despite their wide applications in molecule and protein modeling. Fig. 2 describes the research scopes of Sci-LLMs in this survey.
This survey’s distinct boundaries set it apart from other reviews of LLMs and computational modeling for molecules, proteins, and genomes. In contrast to those primarily centered on natural languages (Zhao et al., 2023; Yang et al., 2023), our emphasis leans more towards scientific languages. Unlike surveys concentrating solely on molecule (Du et al., 2022; Xia et al., 2023), protein (Bepler and Berger, 2021; Hu et al., 2022; Tran et al., 2023; Unsal et al., 2022), or genome data (Consens et al., 2023), we aim to provide a comprehensive view of language models for chemical and biological research. Furthermore, we delve into multimodal LLMs, exploring the interactions between texts and molecule/protein/genome languages. To the best of our knowledge, this nuanced exploration has not been covered in previous surveys. The contributions of the survey can be summarized as follows:
  • We offer a comprehensive review of language modeling within scientific domains, encompassing textual, molecular, protein, and genomic languages, emphasizing domain-specific knowledge.
  • We provide a detailed summary of existing Sci-LLMs, covering model architectures, capabilities, training data, evaluation benchmarks, and assessment criteria. We also show the evolutionary tree of Sci-LLMs in Fig. 3.
  • We enumerate available resources for Sci-LLMs, open-source and maintain the related materials at https://github.com/HICAI-ZJU/Scientific-LLM-Survey, thereby facilitating accessibility for newcomers to the field.
  • To our best knowledge, this survey represents the first comprehensive overview of multimodal Sci-LLMs designed to explore the interaction between various scientific languages.
Figure 2. Research scopes of Scientific Large Language Models (Sci-LLMs) in this survey. We focus on scientific languages (i.e., textual, molecular, protein and genomic languages), as well as their combination (i.e., multimodal language), within the realm of biochemical science.
The remainder of this survey is organized as follows: Sec.2 introduces the background of LLMs and formulates relevant concepts. Sec.3, Sec.4, Sec.5, Sec.6 and Sec.7 present Sci-LLMs, including textual, molecular, protein, genomic and multimodal LLMs, respectively. Finally, we analyze the limitations of existing models, pinpoint potential research directions and conclude this survey in Sec.8.
Figure 3. An evolutionary tree of Sci-LLMs, which consists of five main branches corresponding to the research scopes in this survey. Due to the extensive number of Sci-LLMs, it is not feasible to include all of them in this figure, despite their exceptional quality. For detailed information on the featured models, please refer to Table 1, 3, 5, 7, and 9. Additionally, we will share the source file of this figure and encourage readers to make incremental updates at https://kdocs.cn/l/cbRA94QwMhmn.
2. Background
2.1. Formulation of Scientific Languages
Language is a sophisticated system of communication, typically with well-defined vocabulary and grammatical rules (Evans and Levinson, 2009). It enables the transmission of intricate semantic information, serving as a remarkable feature that distinguishes humans from animals (Le Guin, 2004). One typical language system is the human language, or called natural language, such as English, Chinese and Spanish. In addition to these, there exist non-natural language systems, such as programming languages like Java and Python, and chemical molecule languages like SMILES. Each language system possesses its unique set of vocabulary and grammar, often entirely disparate from others. Consequently, proficiency in one language system does not necessarily equate to proficiency in another (AI4Science and Quantum, 2023). In this survey, we delve into the realm of scientific languages, concentrating not only on domain-specific languages like those of molecules, proteins, and genomes but also on the textual languages that are central to various scientific domains.
Molecular Language. The chemical molecule language system encompasses specialized languages utilized to represent molecular structures and compositions in chemistry. A prominent example in chemical molecule language systems is SMILES (Simplified Molecular Input Line Entry System) (Weininger, 1988). It describes molecular structures by encoding atoms, bonds, and connectivity patterns using ASCII strings. For instance, the SMILES notation for Ethanol is “CCO”. Each letter represents an atom (e.g., ‘C’ for carbon, ‘O’ for oxygen), while special characters denote different bond types (e.g., ‘-’, ‘=’, ‘#’). SELFIES (Simplified Molecular Input Line Entry System) (Krenn et al., 2020) is a recent molecular language that utilizes a syntax designed with specific rules to represent molecular structures. It employs another set of predefined symbols to encode molecular information. For instance, the SELFIES representation of ethanol is “[C][C][O]”. The grammar dictates the rules and syntax for constructing valid representations of molecular structures, and governs how atoms, bonds, and other chemical entities are arranged and connected within the encoded strings. InChI (International Chemical Identifier) (Heller et al., 2013) is another identifier for chemical substances that provides a machine-readable description of the substance. InChI comprises layers of information that can represent molecular connectivity, hydrogen position, isotope composition, stereochemistry, and electronic charge. The InChI algorithm converts input structural information of molecules into a unique identifier through normalization to remove redundant information, canonicalization to generate a unique number label for each atom, and serialization to produce a string of characters. The International Union of Pure and Applied Chemistry (IUPAC) defines a set of rules that converts chemical structures to natural languages (Favre and Powell, 2013).
Protein Language. The protein language system is a specialized framework used to represent the primary structures of proteins. It offers a method to encode amino acid sequences, which are the building blocks of proteins, into a structured format that can be interpreted and analyzed computationally. The vocabulary in the protein language system consists of symbols representing the 20 standard amino acids. Each amino acid is denoted by a specific one-letter or three-letter code (GDR et al., 1984). For instance, ‘A’ represents alanine, ‘R’ represents arginine, and ‘L’ represents leucine in the one-letter code, while ‘ALA’ represents alanine, ‘ARG’ represents arginine, and ‘LEU’ represents leucine in the three-letter code. In terms of grammatical rules, the protein language system does not have a traditional grammar like natural languages. The protein language system enables scientists to describe and analyze protein structures and functions computationally. It allows for the systematic representation of amino acid sequences, aiding in the understanding of protein characteristics, interactions, and biological roles in various biological processes.
Genomic Language. The genomic language system refers to the code used to represent genetic information in living organisms, encompassing both DNA (DeoxyriboNucleic Acid) and RNA (RiboNucleic Acid) (Soukup, 2001). These systems serve as the fundamental languages for encoding genetic instructions and hereditary traits. The vocabulary in the genomic language system includes nucleotide base pairs. In DNA, these are Adenine (A), Thymine (T), Cytosine (C), and Guanine (G). In RNA, thymine is replaced by Uracil (U). Combinations of these bases form the genetic code, representing sequences in genes. Similar to the protein language system, the genomic language does not have a clear set of grammatical rules to organize the tokens in the vocabulary. The genomic language system provides a standardized way to encode and interpret genetic information. It allows scientists to study DNA and RNA sequences, decode genetic instructions, and understand the processes of inheritance, gene expression, and molecular biology.
It is noteworthy that, in addition to the aforementioned sequence-based languages (one-dimensional), alternative forms can be employed to represent molecules, proteins and genomes, such as graphs (two-dimensional) and structures (three-dimensional). However, considering our focus on language modeling with sequences (i.e., LLMs), we have intentionally excluded the models that are solely based on other modalities in this survey.
Figure 4. Illustration of molecular, protein and genomic languages. Molecular languages include SMILEs, SELFIES and InChl sequences, and 2D topology and 3D geometry structures. Protein languages consist of primary structure (i.e., amino acid sequence), secondary, tertiary, and quaternary structures (3D). Genomic languages include DNA and RNA sequences/structures. This survey focuses solely on sequence modeling of molecular, protein and genomic languages.
2.2. Taxonomy of Model Architectures
The taxonomy of model architectures in LLMs can be broadly categorized into three types: encoder-only, decoder-only, and encoder-decoder. Each of these architectures serves different purposes and applies to different kinds of tasks.
Encoder-only. Encoder-only models are a type of neural network architecture that solely consists of an encoder part without a corresponding decoder. These models are designed to process input data and generate a fixed-size representation or encoding of that data. The encoder processes the input sequence and compresses the information into a latent space or encoding which can be used for various downstream tasks such as classification, clustering, or any other task requiring an understanding of the input data. A common example of an encoder-only model is the BERT (Devlin et al., 2019) (Bidirectional Encoder Representations from Transformers) model which is used in natural language processing tasks to generate meaningful representations of text data. In the context of protein, encoder-only models can be employed to learn representations of protein sequences which can then be used for various bioinformatics tasks.
Decoder-only. Decoder-only models are neural network architectures that consist solely of a decoder part without a corresponding encoder. These models are typically used for generating sequences or other forms of output data based on given inputs. Unlike encoder-only models, which aim to create a compact representation of the input data, decoder-only models focus on generating or expanding information, often based on a given seed input or condition. A well-known example of a decoder-only model is the Generative Pre-trained Transformer (GPT (OpenAI, 2022)) series, which is designed to generate text sequences based on provided prompts or initial text fragments. In the field of genomics, decoder-only models could be used for tasks like sequence generation, prediction of missing or subsequent data based on given sequence fragments, or other generative tasks.
Encoder-decoder. Encoder-decoder models are a class of neural network architectures that are particularly powerful in tasks involving sequence-to-sequence transformations. These models consist of two primary components: an encoder and a decoder. The encoder’s function is to process the input data and compress it into a context-rich representation, often referred to as the context vector or encoded state. The decoder, on the other hand, takes the encoded representation and generates the output sequence. A distinguished example is the Text-to-Text Transfer Transformer (T5 (Raffel et al., 2020)), which proposes a unified framework that converts all of natural language processing tasks into a sequence-to-sequence format. The separation of the encoding and decoding processes enables the model to specialize in understanding and generating sequences, making it ideal for various downstream tasks, such as molecule captioning, editing and summarization.
Figure 5. Illustration of common architectures of Sci-LLMs, including (a) encoder-only, (b) decoder-only, and (c) encoder&decoder-based models, with (d) representing the tokens of scientific languages.
2.3. Pre-training and Fine-tuning
The training of LLMs, particularly those designed for the scientific domain, typically involves a two-stage process: pre-training and fine-tuning.
Pre-training. During the pre-training phase, the model is exposed to a vast and diverse corpus of natural and scientific language. For instance, in the natural language-based Sci-LLMs, this corpus may be heavily weighted toward scientific literature, including research papers, academic journals, and technical documents, to imbue the model with a strong grasp of scientific terminology and concepts. The objective function of pre-training is dependent on the chosen model architectures. For example, Sci-LLMs based on BERT (Devlin et al., 2019) often employ masked language modeling, where the goal is to predict the original identity of these masked tokens. This approach allows Sci-LLMs to learn a bidirectional representation of language by considering the contextual information from both sides of the masked token. For Sci-LLMs based on GPT (OpenAI, 2022), the objective can be causal (or autoregressive) language modeling, which predicts the next token given all the previous tokens.
Fine-tuning. The fine-tuning process is crucial for adapting pre-trained language models to specific tasks, such as function prediction and targeted sequence generation. During this stage, Sci-LLMs are fine-tuned with varying objectives and datasets that contain labels. The choice of objective functions also relies on model architectures and specific downstream tasks at hand. For instance, in the chemical reaction classification task, the fine-tuning dataset would consist of reaction samples annotated with type labels, and the optimization objective is to minimize the cross-entropy between the output and ground-truth. For generation tasks, the fine-tuning datasets typically comprise textual instruction data, consisting of pairs of input instructions and corresponding outputs that guide the model to comprehend specific tasks. The instruction data can be sourced from the benchmarks listed in Table 2, or generated from the scientific corpora as described in DARWIN (Xie et al., 2023). In this context, the objective of instruction fine-tuning could be causal language modeling, which is often used in GPT-based Sci-LLMs.
It should be noted that the pipeline of training in Sci-LLMs (including textual, molecular, protein, genomic and multimodal) is consistent with that of general LLMs; therefore, we will not cover the training part (except for training datasets) in their respective chapters.
2.4. Notions and Terms
In this subsection, we clarify the key concepts that will be utilized throughout our discussion. We will not define the terminology like Transformer and pretraining as we assume the readers have a preliminary knowledge of LLMs.
Language. Language is a sophisticated system of communication with well-defined vocabulary and grammatical rules. It includes natural languages like English, used by humans, and non-natural languages such as programming languages crafted by computer scientists, SMILES notation used in chemistry, and amino acid sequences in biology. Scientific Language. Scientific Language is a specialized subset of language, crafted specifically to articulate scientific knowledge. In the context of this survey, we define scientific languages in two broad categories: one, as textual language that is focused on various science domains; and two, as domain-specific languages, which include the languages of molecules, proteins, and genomes. This bifurcation helps in understanding both the general and specialized aspects of language used in scientific discourse and research. Large Language Models (LLMs). LLMs are large-scale neural networks with millions or even billions of trainable parameters, typically based on the Transformer architectures. These models are pre-trained with vast amounts of language data, enabling them to exhibit proficiency in comprehending, generating, and translating natural language.
Scientific Large Language Models (Sci-LLMs). Sci-LLMs are a collective term of LLMs designed to understand, interpret, and generate scientific languages. In this survey, Sci-LLMs include textual, molecular, protein, genomic, and multimodal scientific LLMs. Textual Scientific Large Language Models (Text-Sci-LLMs). Text-Sci-LLMs are a category of LLMs that are specifically trained on vast amounts of textual scientific data. These models excel in understanding, generating, and interacting with human language in its written form. Molecular Large Language Models (Mol-LLMs). Mol-LLMs are specialized LLMs trained on molecular data, allowing them to understand and predict chemical properties and behaviors of molecules. This distinctive expertise renders them invaluable tools in drug discovery, materials science, and understanding complex chemical interactions. Protein Large Language Models (Prot-LLMs). Prot-LLMs are trained specifically on protein-related data, including amino acid sequences, protein folding patterns, and other biological data relevant to proteins. Consequently, they possess the capability to accurately predict protein structures, functions, and interactions. Genomic Large Language Models (Gene-LLMs). Gene-LLMs, with a primary focus on genomic data, are trained to understand and predict aspects of genetics and genomics. They can be used to analyze DNA sequences, understand genetic variations, and assist in genetic research endeavors such as identifying genetic markers for diseases or understanding evolutionary biology. Multimodal Scientific Large Language Models (MM-Sci-LLMs). MM-Sci-LLMs are the most advanced and versatile models, capable of processing and integrating multiple types of scientific data. They can handle text, molecules, proteins, and more, making them suitable for complex scientific research that spans different data types. This capability makes them invaluable in interdisciplinary research areas, where they can synthesize knowledge from various scientific domains to provide comprehensive insights.
3. Textual Scientific Large Language Models
In this section, we aim to explore and delve into scientific large language models specifically trained using textual corpora (i.e., Text-Sci-LLMs), particularly emphasizing their acquisition of chemical and biological knowledge. We will briefly review the existing Text-Sci-LLMs and examine their capabilities, the employed datasets, as well as the evaluation approach. The overview of this section is shown in Figure 6.
{forest}
for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw, rounded corners,align=left, minimum width=2.5em, inner xsep=4pt, inner ysep=1pt, , where level=1text width=5em,fill=blue!10, where level=2text width=5em,font=,fill=pink!30, where level=3font=,yshift=0.26pt,fill=yellow!20, [Text-Sci-LLMs, fill=green!10 [Models
(Sec. 3.1),text width=4em [Biological, text width=4em [Encoder-only: BioBERT (Lee et al., 2020)   BlueBERT (Peng et al., 2020)  PubMedBERT (Gu et al., 2021)  BioMegatron (Shin et al., 2020) ] [Decoder-only: BioGPT (Luo et al., 2022b)   BioMedGPT-LM (Luo et al., 2023c) ] ] [Chemical, text width=4em [Encoder-only: ChemBERT (Guo et al., 2021) MatSciBERT (Gupta et al., 2022) MaterialsBERT (Shetty et al., 2023) ] ] [Comprehen., text width=4em [ Encoder-only: SciBERT (Beltagy et al., 2019) ScholarBERT (Hong et al., 2023) ] [Decoder-only: DARWIN-Base  (Xie et al., 2023) ] ] ] [Datasets
(Sec. 3.2),text width=4em [Pre-training, text width=4em [ PubMed   PubChem   bioRxiv   arXiv   WIPO  etc. ] ] [Benchmarks, text width=4em [PubMedQA (Jin et al., 2019)   MMLU (Hendrycks et al., 2021)   C-Eval (Huang et al., 2023)   AGIEval (Zhong et al., 2023) ScienceQA (Lu et al., 2022)
Bioinfo-Bench-QA (Chen and Deng, 2023)   SciQ (Johannes Welbl, 2017)  Xiezhi (Gu et al., 2023)   ARC (Clark et al., 2018)   SciEval (Sun et al., 2023) ] ] ] [Evaluation
(Sec. 3.3),text width=4em [Capabilities, text width=4em [Pre-college level: Understand basic knowledge ] [College level: Master specialized knowledge ] [Post-college level: Propose innovative knowledge ] ] [Criteria, text width=4em [ Discriminative tasks: Accuracy  Precision  Recall  F1-Score
] [ Generative tasks: BLUE (Papineni et al., 2002)  ROUGE (Lin, 2004)  BERT-Score (Zhang et al., 2019) ] [ Human or GPT-4 (OpenAI, 2023) ] ] ] [Summary
(Sec. 3.4),text width=4em ] ]
Figure 6. Chapter overview of Text-Sci-LLMs.
3.1. Models
To provide a clear overview of Text-Sci-LLMs, we divide them into three distinct categories: biological, chemical, and comprehensive, based on the specific scientific domains they focus on. Table 1 shows the summary of Text-Sci-LLMs.
Biological Domain. Large language models trained on extensive biological corpora, such as BERT (Devlin et al., 2019) and its variations with the encoder-only LLM architecture, have demonstrated significant potential in Natural Language Processing (NLP) tasks within biology. BioBert (Lee et al., 2019), BioMegatron (Shin et al., 2020), PubMedBERT (Gu et al., 2021), BioM-BERT (Alrowili and Shanker, 2021) and LinkBERT (Yasunaga et al., 2022) initially trained on broad corpora like Wikipedia and textbooks and then fine-tuned on specific biological NLP tasks, showcase their significant improvements in various downstream tasks such as biological terminology understanding, named entity recognition, text similarity, and relation extraction. BlueBERT (Peng et al., 2020) presented a thorough empirical study on the efficacy of multi-task learning on BERT-based models for biomedical texts. Since the encoder-only models lack the capability of textual information generation, GPT and its variants (Radford et al., 2018, 2019a; Brown et al., 2020) with a decoder-only architecture have emerged as dominant players in the field of biological NLP. BioGPT (Luo et al., 2022b), an extension of GPT-2 (Radford et al., 2019b), has been extensively pretrained on biomedical literature, showcasing remarkable performance in relation extraction and question answering. Moreover, it demonstrates the ability to generate coherent and fluent descriptions within the biomedical context. BioMedGPT-LM (Luo et al., 2023c), which is incrementally pretrained on LLaMA2 (Touvron et al., 2023a), enabling the comprehensive understanding of various biological modalities and aligning them with natural language.
Table 1. Summary of Text-Sci-LLMs
Domain Model Time #Parameters Base model Pretraining dataset Open-source
Biology BioBERT (Lee et al., 2020) 2019.05 117M BERT PubMed, PMC
BlueBERT (Peng et al., 2020) 2019.07 117M BERT PubMed
BioMegatron (Shin et al., 2020) 2020.10 345M-1.2B BERT PubMed, PMC
PubMedBERT (Gu et al., 2021) 2020.10 117M BERT PubMed $\times$
BioM-BERT(Alrowili and Shanker, 2021) 2021.06 235M BERT PubMed, PMC
BioLinkBERT(Yasunaga et al., 2022) 2022.03 110M, 340M BERT PubMed
BioGPT (Luo et al., 2022b) 2023.03 347M GPT PubMed
BioMedGPT-LM (Luo et al., 2023c) 2023.08 7B LLaMA PMC, arXiv, WIPO
Chemistry ChemBERT (Guo et al., 2021) 2021.06 120M BERT Chemical journals
MatSciBERT  (Gupta et al., 2022) 2021.09 117M BERT Elsevier journals
MaterialsBERT  (Shetty et al., 2023) 2022.09 - BERT Material journals
Comprehensive SciBERT (Beltagy et al., 2019) 2019.09 117M BERT Semantic Scholar
ScholarBERT (Hong et al., 2023) 2023.05 340M, 770M BERT Wiki, Books, etc.
DARWIN-Base (Xie et al., 2023) 2023.08 7B LLaMA SciQ, Web of Science
Chemical Domain. In the realm of chemistry, natural language is commonly used to express various attributes and discoveries related to compounds and chemical reactions, making LLMs well-suited for comprehending and reasoning about chemical information (Bran and Schwaller, 2023). Built upon the BERT architecture, ChemBERT (Guo et al., 2021) was pretrained on an extensive dataset comprising 200,000 chemical journal articles. This model is tailored for automated chemical reaction extraction from chemical literatures, employing Masked Language Modeling (MLM) for pretraining and fine-tuning for chemical NLP tasks. MatSciBERT (Gupta et al., 2022) and MaterialsBERT (Shetty et al., 2023) are materials science-specific models, trained on a large corpus of material literature, and show cutting-edge performance in extracting material property information. It is worth noting that many works utilized ChatGPT (Ouyang et al., 2022) or GPT-4 (OpenAI, 2023) to extract valuable chemical information from literature, such as synthesis conditions and compound properties. For example, Zheng et al. (Zheng et al., 2023b) demonstrated the potential of ChatGPT for text mining and prediction of MOF Synthesis. ChemCrow (Bran et al., 2023) augmented the performance of ChatGPT in chemistry by incorporating 13 expert-designed tools, specifically tailored for accomplishing tasks related to organic synthesis, drug discovery, and materials design.
Comprehensive Domain. This part introduces general Text-Sci-LLMs, which span diverse domains including biology, chemistry, and other scientific disciplines. SciBERT (Beltagy et al., 2019) is built upon the BERT architecture and utilizes unsupervised pretraining on a vast corpus of scientific publications to improve performance on downstream scientific NLP tasks. ScholarBERT (Hong et al., 2023) is a series of BERT-based scientific language models trained on a substantial and diverse scientific corpus spanning domains including life sciences, biomedicine and physical sciences. DARWIN-Base (Xie et al., 2023), based on a decoder-only architecture, is trained on a massive corpus from various scientific texts and knowledge bases. Notably, it was fine-tuned with instruction data from the SciQ (Welbl et al., 2017) dataset and additional question-answer pairs, exhibiting proficiency in tasks that require extensive scientific knowledge. We emphasize that, although the general LLMs such as ChatGPT (Ouyang et al., 2022), GPT4 (OpenAI, 2023) and Gemini (Team, 2023) encompass a wealth of scientific knowledge, they are not tailored for Sci-LLMs, and therefore do not fall under the category of general Text-Sci-LLMs.
3.2. Datasets
According to the availability of annotations, datasets can be categorized into two types: unlabeled data and labeled data. The former is often used for self-supervised pre-training, while the latter is utilized for supervised fine-tuning or model evaluation.
Pre-training Datasets. In Text-Sci-LLMs, the pre-training corpora are mainly sourced from various scientific articles. We list several widely-used data sources as follows.
  • PubMed [^2] https://pubmed.ncbi.nlm.nih.gov/is a free and open-access database in the field of biomedical and life sciences, which contains more than 32 million citations and abstracts of biomedical literature. It is managed by the National Center for Biotechnology Information (NCBI) at the U.S. National Library of Medicine.
  • PubMed Central (PMC) [^3] https://www.ncbi.nlm.nih.gov/pmc/ is a free full-text archive of biomedical and life sciences journals, which has an extensive collection of over 8 million articles. It can be seen as a complementary resource to PubMed, focusing on the storage and free accessibility of full-text papers.
  • PubChem [^4] https://pubchem.ncbi.nlm.nih.gov/ is a database of chemical molecules and their activities against biological assays. This database includes information on chemical structures, identifiers, chemical and physical properties, biological activities, patents, health, safety, toxicity data, and more.
  • bioRxiv [^5] https://www.biorxiv.org/ is an open-access preprint repository for biological sciences, which contains over 2 million full-text papers. This repository covers a wide range of topics in biology, including genetics, neuroscience, bioinformatics, and ecology.
  • arXiv [^6] https://arxiv.org/ is an online repository that provides open access to over 2 million scholarly articles in various disciplines including physics, mathematics, computer science, quantitative biology, quantitative finance, statistics, electrical engineering and systems science, and economics.
  • Semantic Scholar [^7] https://www.semanticscholar.org/ is an advanced, free, AI-powered research tool for scientific literature that provides access to 200 million academic papers sourced from publisher partnerships, data providers, and web crawls.
  • World Intellectual Property Organization (WIPO) [^8] https://www.wipo.int/ serves as a global forum for intellectual property services, policy, information, and cooperation. It maintains an extensive database of patent information that can be accessed by researchers and inventors around the world.
  • Web of Science [^9] https://www.webofscience.com is a paid-access platform that offers access to diverse databases containing reference and citation data from academic journals, conference proceedings, and other scholarly documents across multiple academic disciplines.
Benchmarks. We curate a collection of benchmarks for Text-Sci-LLMs, as presented in Table 2. These benchmarks include multiple subsets that can be utilized to evaluate LLMs’ abilities in grasping scientific knowledge at different levels within the biological, chemical and comprehensive domains.
  • MMLU (Massive Multitask Language Understanding) (Hendrycks et al., 2021). MMLU offers a detailed and challenging benchmark that tests the comprehension and problem-solving capabilities of LLMs across a wide spectrum of tasks and subjects. Specifically, it covers 57 subjects, including STEM (Science, Technology, Engineering, Mathematics), humanities, social sciences, and other fields. Moreover, it ranges in difficulty from elementary to advanced professional levels. We can single out the subjects of biology and chemistry at the high school and college levels to evaluate Text-Sci-LLMs.
  • C-Eval (Huang et al., 2023). C-Eval consists of 13,948 multi-choice questions spanning 52 diverse disciplines and four difficulty levels: middle school, high school, college, and professional level. In terms of scientific domains of interest, C-Eval covers the knowledge of chemistry and biology from middle school to undergraduate levels.
  • AGIEval (Zhong et al., 2023). AGIEval is derived from rigorous and authoritative human-centric examinations, such as college entrance exams and professional certification tests. These exams ensure a robust and standardized evaluation of language models. AGI encompasses knowledge at both the pre-college and college levels. However, in the specific domains of biology and chemistry, its scope is limited to pre-college level knowledge.
  • ScienceQA (Lu et al., 2022). ScienceQA is a benchmark designed to evaluate the ability and interpretability of LLMs in answering multimodal multiple-choice science questions. It consists of ~21k questions covering diverse science topics, including biology and chemistry, as well as other scientific fields.
  • Xiezhi (Gu et al., 2023). Xiezhi is a comprehensive evaluation suite for LLMs. It consists of 249,587 multi-choice questions spanning 516 diverse disciplines, which is sourced from a variety of examinations, including those conducted at the elementary school level, middle school entrance evaluations, college entrance assessments, undergraduate examinations, graduate entrance tests, and adult education exams. Xiezhi encompasses a wide range of scientific subjects, including physics, chemistry, geography, mathematics, physics, biology, medicine, providing over 30k data.
  • SciEval (Sun et al., 2023). SciEval consists of 18,000 challenging scientific questions across chemistry, physics, and biology, assessing LLMs’ capabilities in basic knowledge, knowledge application, scientific calculation, and research ability. It combines objective and subjective questions and dynamically generates data based on basic scientific principles.
  • Bioinfo-Bench-QA (Chen and Deng, 2023). Bioinfo-Bench-QA focuses on evaluating the advanced abilities of foundational models within the context of bioinformatics. It includes 1,500 multiple-choice questions generated by ChatGPT using authoritative bioinformatics literature.
  • BLURB (Gu et al., 2021) focuses on PubMed-based biomedical applications, encompassing a comprehensive compilation of publicly available datasets for various biomedical NLP tasks. These tasks span diverse domains such as named entity recognition, evidence-based medical information extraction, relation extraction, sentence similarity, document classification, and question answering.
  • PubMedQA (Jin et al., 2019). PubMedQA is a biomedical Question-Answering (QA) dataset collected from PubMed, which has 1k expert-annotated, 61.2k unlabeled and 211.3k artificially generated QA instances.
  • ARC (Clark et al., 2018). ARC is comprised of 7,787 questions in the field of natural science, specifically questions that were created for standardized tests. ARC has been divided into two sets: the Challenge Set (consisting of 2,590 questions) and the Easy Set (consisting of 5,197 questions), corresponding to the proposed pre-college and college levels, respectively.
  • SciQ (Johannes Welbl, 2017). SciQ comprises a collection of 13,679 science exam questions on various subjects such as Chemistry, Biology, and more. These questions are presented in a multiple-choice format, offering four answer options for each question. To enhance understanding and knowledge, the majority of the questions are accompanied by an additional paragraph that provides supporting evidence for the correct answer.
Table 2. Summary of the benchmarks for Text-Sci-LLMs
Dataset Time Subset #Item Domain Type Capability Language
MMLU (Hendrycks et al., 2021) 2020.09 High-school-biology 344 Biology Multiple choice Pre-college English
High-school-chemistry 227 Chemistry Pre-college
College-biology 162 Biology College
College-chemistry 110 Chemistry College
C-Eval (Huang et al., 2023) 2023.05 Mid-school-biology 218 Biology Multiple choice Pre-college Chinese
Mid-school-chemistry 210 Chemistry Pre-college
High-school-biology 199 Biology Pre-college
High-school-chemistry 196 Chemistry Pre-college
College-chemistry 253 Chemistry College
AGIEval (Zhong et al., 2023) 2023.04 Gaokao-biology 210 Biology Multiple choice Pre-college Chinese
Gaokao-chemistry 207 Chemistry
ScienceQA (Lu et al., 2022) 2022.09 Natural-science-biology 4098 Biology Multiple choice / QA Pre-college English
Natural-science-chemistry 1194 Chemistry
XieZhi (Gu et al., 2023) 2023.06 Science-biology 2831 Biology Multiple choice Mixed Both
Science-chemistry 399 Chemistry
SciEval (Sun et al., 2023) 2023.08 Basic-biology 2142 Biology Multiple choice / QA Mixed English
Knowledge-biology 1369 Biology
Calculation-biology 299 Biology
Research-biology 995 Biology
Basic-hemistry 2909 Chemistry
Knowledge-chemistry 1700 Chemistry
Calculation-chemistry 3396 Chemistry
Bioinfo-Bench-QA (Chen and Deng, 2023) 2023.10 - 150 Biology Multiple choice Post-college English
BLURB (Gu et al., 2021) 2020.07 - 648k Biology Multiple NLP tasks Mixed
PubMedQA (Jin et al., 2019) 2019.09 - 273.2k Biology True or false College
ARC (Clark et al., 2018) 2018.03 - 7.78k Natural Science Multiple choice Pre-college
SciQ (Johannes Welbl, 2017) 2017.07 - 13.7k Natural Science Multiple choice Mixed
3.3. Evaluation
Capabilities. The evaluation of LLMs is often structured around Bloom’s taxonomy (Krathwohl, 2002; Forehand, 2010), which outlines six cognitive levels of educational learning objectives: Remember, Understand, Apply, Analyze, Evaluate and Create. More recently, SciEval (Sun et al., 2023) has suggested assessing scientific LLMs across four dimensions: basic knowledge, knowledge application, scientific calculation, and research ability. Each dimension aligns with one or more cognitive levels in Bloom’s taxonomy. While these frameworks offer cognitive and research perspectives, our focus is to gauge the depth of knowledge grasped by LLMs, ideally comparable to human comprehension. Therefore, this survey adopts an approach (named KnowEval) that categorizes the capabilities of Text-Sci-LLMs based on the complexity of scientific knowledge, differentiating between Pre-college, College, and Post-college levels, as illustrated in Fig. 7.
  • Pre-college Level. Scientific knowledge at the pre-college level primarily focuses on fundamental concepts, principles, and theoretical frameworks, serving as a foundation for further comprehensive study and research in scientific disciplines. This predominantly textual knowledge is readily accessible across various corpora, making it easily comprehensible for LLMs. We refer to the understanding of basic knowledge as the pre-college level of LLMs, which corresponds to the Remember and Understand levels in Bloom’s taxonomy and the basic knowledge in SciEval. Some evaluation benchmarks such as MMLU (Hendrycks et al., 2021) and C-Eval (Huang et al., 2023) involve assessing pre-college level’s knowledge.
  • College Level. Scientific knowledge at the college level is typically more specialized, abstract, and necessitates logical reasoning and proof rather than mere memorization. It undergoes frequent updates aligned with scientific and industrial advancements, posing a challenge for LLMs to grasp and adapt to these continual changes. The master of specialized knowledge is designated as the college level of LLMs, which typically aligns with the Apply and Analyze levels in Bloom’s taxonomy, as well as the knowledge application and the scientific calculation in SciEval. Most evaluation benchmarks such as PubMedQA (Jin et al., 2019) and SciQ (Johannes Welbl, 2017) focus on this level.
  • Post-college Level. At this level, LLMs need to not only master up-to-date knowledge but also create innovative ideas. The ability to answer questions goes beyond multiple-choice, true/false, and short-answer formats. It also requires the capability to address generative questions, such as summarizing technological advancements and designing novel experimental schemes. This post-college level aligns with the Evaluate and Create levels in Bloom’s taxonomy and the research ability in SciEval. A subset in the SciEval dataset (Sun et al., 2023) is one of the few benchmarks that can reflect such capabilities of LLMs.
Figure 7. Comparsion of evaluation principle in (a) SciEval, (b) Bloom, and (c) KnowEval (this survey).
Evaluation Metric. To quantitatively measure the performance of Text-Sci-LLMs, one can utilize the standard metrics commonly employed in the field of machine learning. These include Accuracy, Precision, Recall, and F1-score, which are particularly well-suited for discriminative tasks such as multiple choice and judgment questions. For generative tasks such as question-answering, it is evaluated by measuring the similarity of the LLM’s output and the ground-truth. These similarity metrics include (but not limited to):
  • BLEU (Bilingual Evaluation Understudy) (Papineni et al., 2002) is a popular metric for evaluating machine translation. BLEU gives a score between 0 and 1 (often expressed as a percentage), where higher scores indicate better translation quality. However, It does not account for the semantic meaning of words and can sometimes rate unnatural translations highly if they have a high n-gram overlap with the reference.
  • ROUGE (Recall-Oriented Understudy for Gisting Evaluation) (Lin, 2004) is primarily used to evaluate automatic text summarization and machine translation. It compares the overlap of n-grams, word sequences, and word pairs between the generated text and a set of reference texts. Higher ROUGE scores signify stronger alignment between the automatic summary and the reference summaries.
  • BERT-Score (Zhang et al., 2019) leverages the contextual embeddings from BERT to evaluate the quality of text generation. It compares the similarity of BERT embeddings between the generated candidate and reference texts. BERT-Score addresses some limitations of BLEU and ROUGE, as it considers the context and semantic meaning of words, leading to a potentially more accurate assessment of text generation quality.
When the ground-truth is absent, one could sort to experts or strong agents like GPT-4 (OpenAI, 2023) for assessing LLMs. Human evaluation plays a pivotal role in assessing the quality and accuracy of model-generated results, as it closely aligns with real-world application scenarios and provides comprehensive and precise feedback. However, due to extensive human involvement, this process is time-consuming and labor-intensive. Therefore, there is a growing tendency that employ GPT-4 as a “judge” to evaluate different models (Bubeck et al., 2023). Overall, these metrics are often used complementarily in various NLP tasks to provide a comprehensive evaluation of LLMs.
3.4. Summary
This section focuses on Sci-LLMs trained with textual scientific corpora, particularly those acquiring extensive chemical and biological knowledge. We carefully review the existing Text-Sci-LLMs, their datasets, and evaluation methods. One can conclude that the development of Text-Sci-LLMs is built upon the advancements of general LLMs, with the key difference being the incorporation of more domain-specific corpora during training and a greater emphasis on evaluating the accuracy of scientific knowledge. However, in contrast to the proliferation of LLMs focused on general domains, the number of Text-Sci-LLMs remains limited. Additionally, there are few benchmarks for evaluating LLMs at the post-college level which is a crucial capability for advanced scientific research.
4. Molecular Large Language Models
Large language models have shown great potential in accelerating chemical molecular discovery. In this section, we provide a review of LLMs trained in molecular language (Mol-LLMs), including insights into their model architectures, capabilities, utilized datasets, and evaluation criteria. The overview of this section is shown in Figure 8.
{forest}
for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw, rounded corners,align=left, minimum width=2.5em, inner xsep=4pt, inner ysep=1pt, , where level=1text width=5em,fill=blue!10, where level=2text width=5em,font=,fill=pink!30, where level=3font=,yshift=0.26pt,fill=yellow!20, [Mol-LLMs, fill=green!10 [Models
(Sec. 4.1),text width=4em [Encoder-only, text width= 5.5em [BERT-based: SMILES-BERT (Wang et al., 2019)   MTL-BERT (Zhang et al., 2022b)  rxnfp-BERT (Schwaller et al., 2021)
Mol-BERT (Li and Jiang, 2021)  MolFormer (Ross et al., 2022)   MolRoPE-BERT (Liu et al., 2023d) ] [RoBERTa-based: ChemBERTa (Chithrananda et al., 2020)  MFBERT (Abdel-Aty and Gould, 2022)
SELFormer (Yüksel et al., 2023)   Semi-RoBERTa (Tran and Ekenna, 2023)  ] [Integrating 2D or 3D graphs: GROVER (Rong et al., 2020)  MAT (Maziarka et al., 2020)  MG-BERT (Zhang et al., 2021)
KPGT (Li et al., 2022)  AGBT (Chen et al., 2021)   Molformer (Wu et al., 2023)  Uni-Mol (Zhou et al., 2023a)  ] ] [Decoder-only, text width= 5.5em [GPT-based: MolGPT (Bagal et al., 2021)  SMILESGPT (Adilov, 2021)  cMolGPT (Wang et al., 2023f)   Taiga (Mazuz et al., 2023)  iupacGPT (Cho et al., 2023) ] ] [Encoder-decoder, text width=5.5em [Vanilla Transformer-based:
Molecular Transformer (Schwaller et al., 2019)  SMILES Transformer (Honda et al., 2019)   SCROP (Zheng et al., 2019)
Retrosynthesis Transformer (Karpov et al., 2019)  ChemReactNet (Tetko et al., 2020)   X-MOL (Xue et al., 2020)
Two-way Transformer (Kim et al., 2021b)   RetroSynth-Diversity (Toniato et al., 2023)  GO-PRO (Mann and Venkatasubramanian, 2021)
Transmol (Zhumagambetov et al., 2021)   RetroTRAE (Ucak et al., 2022)  Disconnection aware model (Thakkar et al., 2023) ] [BART-based: Chemformer (Irwin et al., 2022)  BARTSmiles (Chilingaryan et al., 2022)  MOLGEN (Fang et al., 2023c) ] [Transformer+Graph: GET (Mao et al., 2021)  Graph2SMILES (Tu and Coley, 2022) ] [Others: GCT (Kim et al., 2021c) ] ] ] [Datasets
(Sec. 4.2),text width=4em [Pre-training, text width=4em [ZINC (Sterling and Irwin, 2015)  PubChem (Wang et al., 2012)  USPTO (Schneider et al., 2016)  PCQM4M (Hu et al., 2021; Ying et al., 2021)  GEOM (Axelrod and Gomez-Bombarelli, 2022)
MolTLU (Beaini et al., 2023)  ChEMBL (Gaulton et al., 2012)   DrugBank (Wishart et al., 2018)  GDB-17 (Ruddigkeit et al., 2012)  ExCAPE-DB (Sun et al., 2017)  ] ] [Benchmarks, text width=4em [MoleculeNet (Wu et al., 2018)  MARCEL (Zhu et al., 2023)  GuacaMol (Brown et al., 2019)   MOSES (Polykovskiy et al., 2020)  ADMETlab 2.0  (Xiong et al., 2021) ] ] ] [Evaluation
[…]
(Sec. 4.4),text width=4em ] ]
Figure 8. Chapter overview of Mol-LLMs.
4.1. Models
As Transformer (Vaswani et al., 2017) can effectively model sequential data, numerous Transformer-based Mol-LLMs have been proposed for predicting molecular properties and structures. Here we categorize Mol-LLMs into three distinct types based on their specific architectures: encoder-only, decoder-only, and encoder-decoder models, as shown in Table 3.
4.1.1. Encoder-only Models
Encoder-only models focus on understanding and interpreting the input molecules, making them well-suited for tasks that require a deep comprehension of molecular structures and properties. A typical example is SMILES-BERT (Wang et al., 2019) that leverages the BERT architecture to interpret SMILES representations of molecules. It undergoes pre-training on an extensive dataset of unlabeled data through masked SMILES recovery tasks, enabling it to learn nuanced patterns and relationships within molecular structures. Nevertheless, the fine-tuning phase for individual downstream tasks is inefficient. MTL-BERT (Zhang et al., 2022b) mitigates this limitation by employing a multitask learning framework, which substantially increases the accuracy of predictions for multiple molecular properties. The rxnfp-BERT (Schwaller et al., 2021) model draws on the principles of BERT to process and learn from chemical reactions. It is designed to generate reaction fingerprints, which are representations of chemical reactions used for various predictive tasks. However, these models predominantly concentrate on contextual information in molecular sequences while giving less attention to molecular substructures, such as functional groups. Mol-BERT (Li and Jiang, 2021) extracts atomic-level and substructural features centered on the current atom. MoLFormer (Ross et al., 2022) further enhances this approach by integrating relative position encoding and rotational position embedding, thereby facilitating the comprehension of spatial relationships between atoms in a molecule. MolRoPE-BERT (Liu et al., 2023d) augments this method by employing Rotary Position Embedding (RoPE) to efficiently encode positional information in SMILES, aiming to provide a more robust understanding of molecules.
Table 3. Summary of Mol-LLMs
Model Time #Parameters Base model Pre-training Dataset Capability Open-source
Encoder-only SMILES-BERT (Wang et al., 2019) 2019.09 - BERT ZINC Prop. pred.
MAT (Maziarka et al., 2020) 2020.02 - Trans. enc. ZINC-15 Prop. pred.
ChemBERTa (Chithrananda et al., 2020) 2020.10 - RoBERTa PubChem Prop. pred.
GROVER (Rong et al., 2020) 2020.10 48/100M Trans. enc.+GNN ZINC-15, ChEMBL Prop. pred.
rxnfp-BERT (Schwaller et al., 2021) 2021.01 - BERT Pistachio, USPTO React. pred.
MG-BERT (Zhang et al., 2021) 2021.05 - BERT+GNN ChEMBL Prop. pred.
AGBT (Chen et al., 2021) 2021.06 - BERT ChEMBL Prop. pred.
Mol-BERT (Li and Jiang, 2021) 2021.09 - BERT ZINC-15, ChEMBL 27 Prop. pred.
KPGT (Li et al., 2022) 2022.08 100M Trans. enc. ChEMBL 29 Prop. pred.
ChemBERTa-2 (Ahmad et al., 2022) 2022.09 5/46M RoBERTa PubChem Prop. pred.
MFBERT (Abdel-Aty and Gould, 2022) 2022.10 88M RoBERTa GDB-13, ZINC-15, PubChem, ChEMBL, USPTO Prop. pred.
MolFormer (Ross et al., 2022) 2022.12 - Trans. enc. PubChem, ZINC Prop. pred.
MTL-BERT (Zhang et al., 2022b) 2022.12 - BERT ChEMBL Prop. pred.
MolRoPE-BERT (Liu et al., 2023d) 2023.01 - BERT ZINC-15, ChEMBL 27 Prop. pred. $\times$
Molformer (Wu et al., 2023) 2023.01 - Trans. enc. ZINC, PubChem Prop. pred.
Uni-Mol (Zhou et al., 2023a) 2023.02 47.61M Trans. enc. ZINC, ChemBL PDBbind Prop. pred. Conform. gen.
SELFormer (Yüksel et al., 2023) 2023.05 58/87M RoBERTa ChEMBL Prop. pred.
Semi-RoBERTa (Tran and Ekenna, 2023) 2023.07 - RoBERTa PubChem Prop. pred. $\times$
Decoder-only MolGPT (Bagal et al., 2021) 2021.05 6M GPT MOSES, GuacaMol Mol. gen.
SMILES GPT (Adilov, 2021) 2021.09 13.4M GPT PubChem Mol. gen., Prop. pred.
cTransformer (Wang et al., 2022b) 2022.10 - Trans. dec. MOSES Mol. gen. $\times$
iupacGPT (Cho et al., 2023) 2023.05 1500M GPT PubChem Mol. gen., Prop. pred.
cMolGPT (Wang et al., 2023f) 2023.05 - GPT MOSES Mol. gen.
Taiga (Mazuz et al., 2023) 2023.05 - GPT MOSES, ZINC, GDB 13 Mol. gen., Prop. pred.
Encoder-Decoder Molecular- Transformer (Schwaller et al., 2019) 2019.08 12M Transformer USPTO, Pistachio React. pred.
Retrosynthesis- Transformer (Karpov et al., 2019) 2019.09 - Transformer USPTO React. pred.
SMILES- Transformer (Honda et al., 2019) 2019.11 - Transformer ChEMBL 24 Prop. pred.
SCROP (Zheng et al., 2019) 2019.12 - Transformer USPTO React. pred.
ChemReactNet (Tetko et al., 2020) 2020.11 - Transformer USPTO React. pred.
X-MOL (Xue et al., 2020) 2020.12 - Transformer ZINC-15 Prop. pred., React. pred. Mol. optim., Mol. gen.
Two-way- Transformer (Kim et al., 2021b) 2021.01 - Transformer USPTO React. pred.
GO-PRO (Mann and Venkatasubramanian, 2021) 2021.03 5M Transformer USPTO React. pred. $\times$
Transmol (Zhumagambetov et al., 2021) 2021.07 - Transformer MOSES Mol. gen.
RetroTRAE (Ucak et al., 2022) 2021.08 - Transformer USPTO, PubChem, ChEMBL React. pred.
GET (Mao et al., 2021) 2021.10 - Transformer +GNN USPTO React. pred.
Graph2SMILES (Tu and Coley, 2022) 2021.10 - Transformer USPTO React. pred.
GCT (Kim et al., 2021c) 2021.12 - Transformer MOSES Mol. gen.
Chemformer (Irwin et al., 2022) 2022.01 45/230M BART ZINC-15 Prop. pred., React. pred.
BARTSmiles (Chilingaryan et al., 2022) 2022.11 - BART ZINC 20 Prop. pred.
RetroSynth- Diversity (Toniato et al., 2023) 2023.02 12M Transformer USPTO, Pistachio React. pred.
Disconnection aware- model (Thakkar et al., 2023) 2023.07 - Transformer USPTO Pistachio React. pred.
MOLGEN (Fang et al., 2023c) 2023.10 - BART ZINC-15 Mol. gen.
RoBERTa (Liu et al., 2019), building upon the BERT’s framework, introduces dynamic token masking for varied learning and eliminates the next sentence prediction task to better learn long-range dependencies. These enhancements address BERT’s limitations and elevate RoBERTa’s performance across diverse NLP tasks, making it suitable for specialized applications like molecular modeling. A variety of models have been built upon the RoBERTa framework, including chemBERTa (Chithrananda et al., 2020), chemBERTa-2 (Ahmad et al., 2022), MFBERT (Abdel-Aty and Gould, 2022), SELFormer (Yüksel et al., 2023) and semi-RoBERTa (Tran and Ekenna, 2023). MFBERT and semi-RoBERTa both use SMILES sequences as their input, crafting simple molecular representations. MFBERT (Abdel-Aty and Gould, 2022) employs distributed training and a data-driven approach, focusing on generating targeted molecular fingerprints using SMILES sequences. It uses a sentence-piece tokenization model, ensuring robust and accurate molecular representations. Meanwhile, semi-RoBERTa (Tran and Ekenna, 2023), uses SMILES inputs to capitalize on 3D compound structures.
Although SMILES is commonly used for molecular representation, it poses challenges in terms of validity and robustness. To address these issues, SELFIES can be an alternative to SMILES. As a molecular notation system, SELFIES (Krenn et al., 2020) represents organic molecules in an accessible string format with a recursive syntax, enabling self-referencing to enhance both the accuracy and computational efficiency of representations. ChemBERTa (Chithrananda et al., 2020) adopts a broader approach, processing both SMILES and SELFIES inputs. This model delves into the effects of size, tokenizer types, and string representations, employing Byte-Pair Encoding (BPE) to manage these varied inputs. ChemBERTa-2 (Ahmad et al., 2022) further integrates multi-task regression and larger datasets, thereby augmenting its proficiency in molecular property analysis.
The above BERT-based models integrate a vast amount of unlabeled molecular data through a self-supervised learning strategy, but they are based on the sequential format of molecule SMILES or SELFIES, overlooking crucial graph structures and three-dimensional (3D) chemical information. The oversight of graph structures can be effectively addressed by integrating graphs with Transformers, such as GROVER (Rong et al., 2020), MAT (Maziarka et al., 2020), MG-BERT (Zhang et al., 2021), and KPGT (Li et al., 2022). For example, MG-BERT (Zhang et al., 2021) combines GNN’s local message passing with BERT’s architecture, enhancing molecular feature extraction and utilizing a unique pretraining strategy involving atom masking. Certainly, GNNs alone can deal with molecule graph structures, and representative works include KCL (Fang et al., 2022), KANO (Fang et al., 2023d), GS-Meta (Zhuang et al., 2023b), and iMoLD (Zhuang et al., 2023a). Moreover, several approaches have been employed to model 3D molecules, including AGBT (Chen et al., 2021), Molformer (Wu et al., 2023), and Uni-mol (Zhou et al., 2023a), and DeepSorption (Cui et al., 2023), showcasing the evolution and diversification of computational methods for molecular modeling. For example, DeepSorption leverages a SE(3)-equivalent Transformer to effectively model the 3D structures of molecular crystals, leading to accurate prediction of gas adsorption. Graph and 3D structure-based molecular representation learning is out the scope of this survey and we refer interested readers to the surveys (Guo et al., 2023; Xia et al., 2023).
4.1.2. Decoder-only Models
In recent years, decoder-only architectures like GPT have revolutionized molecular generation in cheminformatics. Mol-LLMs based on the GPT architecture, including MolGPT (Bagal et al., 2021), SMILESGPT (Adilov, 2021), iupaccGPT (Cho et al., 2023), cMolGPT (Wang et al., 2023f), and Taiga (Wang et al., 2022b), primarily use SMILES strings as input, effectively addressing the challenge of navigating the vast chemical space. These models are pivotal in drug discovery and material science, enabling the synthesis of molecules with specific properties.
MolGPT (Bagal et al., 2021), a pioneer in utilizing GPT for molecular generation, uses conditional training for property optimization. It stands out for its efficiency and effectiveness in molecular modeling and drug discovery, showing strong control over multiple properties for accurate generation. Moreover, both SMILESGPT and iupacGPT are based on the GPT-2 architecture. SMILESGPT (Adilov, 2021) is a causative Transformer tailored for drug discovery, training on SMILES notation with BPE, and employing special embedding layers for effective molecular representation learning. iupacGPT (Cho et al., 2023) diverges by using the more readable IUPAC nomenclature, excelling in learning hierarchical and inner-sequence relationships between atoms and chemical groups. The cTransformer (Wang et al., 2022b) and cMolGPT (Wang et al., 2023f) models are specifically developed for the purpose of generating molecules based on given conditions. cTransformer (Wang et al., 2022b) synergizes the Transformer architecture with conditional embeddings, allowing for target-specific molecule generation and fine-tuning with target-specific data. cMolGPT (Wang et al., 2023f) represents an advancement in this domain by generating target-specific compounds by utilizing target embeddings as keys and values within its multi-head attention, adeptly producing molecules active against designated targets. Taiga (Mazuz et al., 2023) exemplifies the combination of GPT and reinforcement learning (RL) methodologies. It employs a two-stage approach, initially treating the problem as a language modeling task with the prediction of the next token in SMILES strings. Afterwards, it utilizes RL to optimize chemical features, notably emphasizing qualities such as a quantitative estimate of QED (Quantitative Estimate of Drug-likeness). This demonstrates the smooth integration of GPT’s language processing abilities with the accuracy of RL approaches.
4.1.3. Encoder-Decoder Models
In the encoder-decoder architecture, encoders convert raw molecules into latent vectors, and decoders reconstruct these latent vectors into functional chemical structures. In the majority of Transformer-based encoder-decoder models, SMILES or SELFIES serve as inputs for the encoder, and the outputs vary across tasks. For example, in the context of chemical reaction prediction, the decoder generates the anticipated outcomes for reactants. The following paragraphs describe Mol-LLMs based on the encoder-decoder architecture in chronological order.
In the field of chemical reaction prediction, Molecular Transformer (Schwaller et al., 2019) is a pioneering Transformer-based model for reaction prediction, which enables the effective handling of complex, long-range sequence interactions. Retrosynthesis Transformer (Karpov et al., 2019) stands out for snapshot learning techniques to improve training efficiency. It utilizes a position encoding matrix to address the issue of distance between string elements, enhancing its ability to process and interpret complex chemical information. SCROP (Zheng et al., 2019) excels in inferring a set of original candidate reactants and incorporates a syntax corrector with position encoding to address the challenges of string element distances. ChemReactNet (Tetko et al., 2020) excels in chemical reaction prediction by employing advanced data augmentation, notably using SMILES augmentation and beam search algorithms, to significantly reduce neural network data memorization and enhance prediction of new sequences. Two-Way Transformer (Kim et al., 2021b) in retrosynthesis features a novel architecture where modules of two Transformers are tied together, sharing the encoder and most decoder components. It employs latent modeling with a multinomial latent variable, significantly diversifying the range of generated reactant candidates. GO-PRO (Mann and Venkatasubramanian, 2021) combines Transformer with a context-free grammar (CFG)-based representation of molecules. It encodes chemical reactions using a grammar-based system, utilizing either a greedy or beam-search decoding strategy for output, resulting in efficient, structurally informed predictions with reduced model complexity.
BART (Lewis et al., 2020), a Transformer-based architecture renowned for its bidirectional encoder and autoregressive decoder, has been effectively applied in molecular design. Chemformer (Irwin et al., 2022), BARTSmiles (Chilingaryan et al., 2022) and MOLGEN (Fang et al., 2023c) are representative models based on the BART architecture. Chemformer (Irwin et al., 2022), a BART language model fine-tuned for sequence-to-sequence and discriminative tasks in cheminformatics, showcases the efficacy of multi-task learning. It employs three innovative self-supervised pre-training techniques, applied to a substantial dataset of unlabeled SMILES strings. BARTSmiles (Chilingaryan et al., 2022), a robust generative masked language model for molecules, leverages extensive self-supervised pre-training on over 1.7 billion molecules. It stands out for its innovative training scale and denoising objectives, significantly boosting performance across multiple molecular tasks. MOLGEN (Fang et al., 2023c) combines domain-agnostic pre-training with a self-feedback paradigm, emphasizing the generation of chemically valid molecules and aligning generative probabilities with real-world chemical preferences to address challenges such as molecular hallucinations.
Additionally, there are several models that use specialized inputs or combine Transformers with distinctive architectures. RetroTRAE (Ucak et al., 2022) not only excels in fragment-based tokenization but also learns changes in atomic environments during chemical reactions. GCT (Kim et al., 2021c) integrates a Transformer-based language model with a conditional variational autoencoder, augmented by a conditional Gaussian latent space. RetroSynth-Diversity (Toniato et al., 2023) and Disconnection Aware model (Thakkar et al., 2023) both employ prompts to guide their retrosynthesis predictions, enhancing diversity and accuracy in different ways. RetroSynth-Diversity (Toniato et al., 2023) leverages cluster token prompts to diversify retrosynthesis predictions, incorporating classification tokens into the molecular language representation. This approach guides the model towards a wider array of disconnection strategies, significantly broadening the exploration of chemical space. The Disconnection Aware model (Thakkar et al., 2023) enhances retrosynthesis pathway predictions by accurately identifying disconnection sites in molecules, using atom-tagged prompts integrated into product SMILES. It significantly boosts reaction class diversity and requires specialized training with inputs containing tagged disconnection sites.
However, these molecular representation learning methods often overlook the properties of substructures, neglecting natural atomic connections and molecular topological information. GET (Mao et al., 2021) and Graph2SMILES (Tu and Coley, 2022) are graph-based approaches designed to address these gaps. GET (Mao et al., 2021) integrates graph-level and sequence-level representations, combining a novel GNN with the Transformer model to enhance retrosynthesis prediction and generate more accurate SMILES outputs. Graph2SMILES (Tu and Coley, 2022) merges Transformers with molecular graphs, which incorporates a global attention encoder for long-distance and intermolecular interactions.
Table 4. Summary of the datasets for Mol-LLMs.
Dataset Time Subset/Version Scale Keywords/Tasks
Pre-training ZINC 2015.10 ZINC-15 (Sterling and Irwin, 2015) 120M Ligand discovery
2020.12 ZINC-20 (Irwin et al., 2020) 140M
2012.07 ZINC-250k(Irwin et al., 2012) 250k
PubChem (Kim et al., 2023) 2023.01 Compound 111.9M Unique chemical structures
Substance 296.9M Chemical entities
BioAssay 1.5M Biological experiments
Bioactivity 296.8M Biological activity data points
Protein 185.1K Tested and identified proteints
Gene 104K Tested and identified genes
Pathway 239K Interactions between chemicals, genes, and proteins
Cell Line 2K Tested cell lines
Taxonomy 112.6K Organisms of proteins/genes
Patent 42.4M Patents with links in PubChem
Data Sources 872 Organizations contributing data to PubChem
USPTO (Lowe, 2012) 2012.10 USPTO MIT 480k US Patents, unique reactions
USPTO-15K 15k
USPTO-full 950k
PCQM4M 2021.10 PCQM4Mv2 (Hu et al., 2021) 3.7M Quantum property, molecular graphs pred.
2021.06 PCQM4M-LSC (Ying et al., 2021) 3.8M Large-Scale
GEOM (Axelrod and Gomez-Bombarelli, 2022) 2022.04 QM9 134K Conformers
AICures drug dataset 317.9K Organic molecules, high-quality conformers
part of MoleculeNet 16.9K Biophysics, physiology, physical chemistry
MolTLU (Beaini et al., 2023) 2023.10 ToyMix 154K Quantum ML, drug discovery, GNN expressivity
LargeMix 5.4M Biological properties
UltraLarge 83M Quantum properties
ChEMBL (Zdrazil et al., 2023) 2023.05 - 22.7M Drug-like bioactive compounds
DrugBank 5.0 (Wishart et al., 2018) 2017.11 - 500K Drug’s mechanisms, interactions, targets
GDB-17 (Ruddigkeit et al., 2012) 2012.10 - 166.4B Organic small molecules
ExCAPE-DB (Sun et al., 2017) 2017.03 - 70M Bioactivity, Chemogenomics
Benchmarks MoleculeNet (Wu et al., 2018) 2017.10 ESOL 1k Physical chemistry prediction
FreeSoly 1K
Lipophilicity 4K
PCBA 437K Biophysics prediction
MUV 92K
HIV 41K
PDBbind 11K
BACE 1K
BBBP 2K Physiology prediction
Tox21 7K
ToxCast 8K
SIDER 1K
ClinTox 1K
MARCEL (Zhu et al., 2023) 2023.09 Drugs-75K 75k Highly flexible molecules
Kraken 1.5k Organophosphorus ligand sterics
EE 872 Chiral catalysts and enantiomers
BDE 5.9k Organometallic catalyst conformations
GuacaMol (Brown et al., 2019) 2019.03 - - Molecule generation
MOSES (Polykovskiy et al., 2020) 2020.12 ZINC Clean Leads 1.9M Molecule generation
ADMETlab 2.0  (Xiong et al., 2021) 2021.04 - 250K Property prediction
4.2. Datasets
Molecule datasets are generally categorized into two types based on the availability of annotations: unlabeled data and labeled data. Each type plays a different role in the model training and evaluation process.
Pre-training Datasets. Unlabeled data is often used for self-supervised learning, where the model learns to understand and represent molecular structures without explicit guidance on specific tasks. Here is a collection of pre-training datasets for Mol-LLMs, as summarized in Table 4.
  • ZINC (Irwin et al., 2012, 2020; Sterling and Irwin, 2015). ZINC serves as a vital repository for researchers, especially in molecular dynamics simulations and virtual screening, offering an extensive array of compounds, including those ready for immediate delivery. ZINC 250K is a curated subset of approximately 250,000 molecules from ZINC, known for its pharmaceutical relevance and moderate size. ZINC-15 is an advanced iteration, offering billions of compounds and enhanced search capabilities for a growing chemical space. ZINC 20 further expands this with an ultralarge-scale chemical database, incorporating novel molecules and sophisticated search technologies to address the challenges of massive, dynamic chemical libraries.
  • PubChem (Wang et al., 2012). PubChem is a comprehensive database that offers free access to information on chemical substances and their biological activities. It contains over 750 million records covering a wide range of data, including chemical structures, identifiers, bioactivity outcomes, genes, proteins, and patents. PubChem is a crucial public repository for biological activity data of small molecules and RNAi reagents, aimed at providing free and easy access to a vast array of data, including over 130 million bioactivity outcomes, 5000 protein targets, and 30,000 gene targets. This extensive database, housing 500,000 assay protocol descriptions, 1.6 million small molecules, 120 chemical probes, and 60,000 RNAi reagents, is integral to molecular modeling and drug discovery.
  • USPTO(Schneider et al., 2016). The USPTO dataset is a significant compilation of chemical reactions extracted from US patents spanning 1976 to 2015. The dataset includes about 1.3 million unique reactions, covering a broad range of more than 200 different reaction types. This dataset is particularly useful for developing and evaluating new methods in reaction role assignment and atom-to-atom mapping, offering a diverse and extensive collection of reactions for cheminformatics research.
  • PCQM4M (Hu et al., 2021; Ying et al., 2021). The PCQM4Mv2 and PCQM4M-LSC datasets are designed for graph-level prediction tasks in quantum chemistry, primarily focusing on predicting the Density Functional Theory (DFT) calculated HOMO-LUMO energy gaps of molecules from their 2D molecular graphs. These datasets provide a robust platform for developing and benchmarking advanced machine learning models in predicting molecular properties. The primary difference between PCQM4Mv2 and PCQM4M-LSC is that PCQM4Mv2 is an enhanced version with more data and features, whereas PCQM4M-LSC is part of the OGB Large-Scale Challenge, emphasizing the development of state-of-the-art graph ML models with a specific scaffold data split and feature set.
  • GEOM (Axelrod and Gomez-Bombarelli, 2022). The Geometric Ensemble of Molecules (GEOM) dataset is a comprehensive collection designed for molecular property prediction and molecular generation. It encompasses 37 million molecular conformations for over 450,000 molecules, including 133,000 species from QM9 and 317,000 species with experimental data related to biophysics, physiology, and physical chemistry. The GEOM dataset is instrumental in the field of molecular modeling, especially for developing models that predict properties from conformer ensembles and for generative models sampling 3D conformations.
  • MolTLU (Beaini et al., 2023). It is a comprehensive suite of datasets for molecular learning, categorized into ToyMix, LargeMix, and UltraLarge. These categories encompass seven distinct datasets, covering nearly 100 million molecules and over 13 billion labels across more than 3,000 tasks. ToyMix combines QM9, TOX21, and ZINC12K, offering 34 labels per molecule for foundational model training in diverse areas such as quantum machine learning and drug discovery. LargeMix merges datasets from quantum chemistry, bio-assays, and transcriptomics, providing data for millions of compounds. UltraLarge, featuring the PM6_83M dataset, stands out for its vast scale, covering 83 million molecules for training 2D-GNNs. These datasets, with their blend of quantum and biological properties, are crucial for advancing molecular modeling.
  • ChEMBL (Gaulton et al., 2012). The ChEMBL database is a vast repository of bioactive molecules with drug-like properties. It includes over 20.3 million bioactivity measurements derived from scientific experiments and literature, along with 2.4 million unique chemical compounds. These data are crucial for drug discovery and pharmacological research, providing a comprehensive resource for scientists studying the effects of molecular compounds on biological systems. This contrasts with datasets like ZINC, which contain virtual molecules that may be synthesizable but have not yet been created. Crucially, ChEMBL’s data serves as an indispensable training dataset for large molecular model training and evaluation, as exemplified by its use in the GuacaMol benchmarking platform.
  • DrugBank 5.0 (Wishart et al., 2018). DrugBank 5.0 includes over 2,358 approved and 4,501 investigational drugs, along with a dramatic increase in drug-drug interactions and pharmacogenomic data. This comprehensive resource offers detailed information on a broad range of drug targets, including proteins, RNA, and DNA, with notable enhancements in fields such as pharmacometabolomics, pharmacotranscriptomics, and pharmacoproteomics. Additionally, DrugBank 5.0 provides extensive data on clinical trials and integrates advanced search features and spectral data, making it an invaluable tool for pharmacological research and the pharmaceutical industry.
  • GDB-17 (Ruddigkeit et al., 2012). GDB-17 (General Database 17) is a comprehensive database containing 166.4 billion organic small molecules. It includes molecules with up to 17 atoms of C, N, O, S, and halogens, thus considerably expanding the available chemical space beyond that of its predecessors, such as GDB-13 (Blum and Reymond, 2009). The database is characterized by a diverse array of nonaromatic heterocycles and various scaffold types, encompassing millions of isomers of known drugs. GDB-17 is especially significant in molecular modeling and drug discovery, providing a broad spectrum of molecular structures for virtual screening and the discovery of novel bioactive compounds.
  • ExCAPE-DB (Sun et al., 2017). ExCAPE-DB, a substantial chemogenomics dataset, amalgamates data from ChEMBL and PubChem, comprising 998,131 unique compounds and 70,850,163 SAR data points across 1,667 targets. It is essential for constructing Molecular Activity Quantification Structure (QSAR) models to predict activities against specific targets and serves as a benchmark for appraising various machine-learning algorithms, particularly those focusing on multi-target learning. Reflecting the pharmaceutical industry’s chemogenomics datasets, ExCAPE-DB includes a significant proportion of inactive compounds, offering a more accurate bioactivity profile. It covers key target families like enzymes, membrane receptors, ion channels, and transcription factors, with most compounds adhering to the Lipinski rule of five, denoting their drug-like characteristics.
Benchmarks. We also curate a collection of benchmarks for Mol-LLMs (in Table 4), which can be effectively utilized for training and evaluating the performance in molecular generation and prediction tasks.
  • MoleculeNet (Wu et al., 2018). MoleculeNet, a benchmark dataset comprising over 700K compounds from multiple public databases, is a cornerstone in the molecular sciences for the development and evaluation of machine learning models. It encompasses four key categories: quantum mechanics, physical chemistry, biophysics, and physiology, each targeting a distinct molecular science aspect. Quantum mechanics datasets (QM7, QM7b, QM8, and QM9) emphasize the electronic properties, molecular geometries, and energy profiles of small molecules, which are crucial for computational chemistry. Physical Chemistry datasets (ESOL, FreeSolv, and Lipophilicity) provide insights into solvation and molecular interaction, impacting drug discovery and material science. Biophysics datasets (PCBA, MUV, HIV, PDBbind, and BACE) offer crucial data on protein-ligand binding affinities and various biophysical interactions. Physiology datasets (BBBP, Tox21, ToxCast, SIDER, and ClinTox) focus on molecular physiological effects, vital for assessing drug safety and environmental toxicity.
  • MARCEL (Zhu et al., 2023). The MoleculAR Conformer Ensemble Learning (MARCEL) benchmark provides a comprehensive platform for evaluating the potential of learning from molecular conformer ensembles. MARCEL marks a significant shift in molecular representation learning by framing it as an ensemble learning task, focusing on diverse molecular conformer structures. This approach is based on the understanding of molecules’ inherent flexibility, characterized by their ability to adopt different conformations due to bond rotations and minor vibrational perturbations. Moreover, MARCEL includes four diverse datasets encompassing a wide range of molecule- and reaction-level properties.
  • GuacaMol (Brown et al., 2019). GuacaMol is an evaluation framework designed for de novo molecular design, which aims to generate molecules with specific property profiles through virtual design-make-test cycles. It offers a suite of standardized benchmarks to measure the models’ fidelity in reproducing the property distribution of training sets, their ability to generate novel molecules, and their efficiency in exploring and exploiting chemical space through a range of single and multi-objective optimization tasks.
  • MOSES (Polykovskiy et al., 2020). Molecular Sets (MOSES) is a benchmark suite for molecular generation. MOSES includes standardized datasets, data preprocessing tools, evaluation metrics, and molecular generation models. The dataset contains ZINC Clean Leads Dataset, which consists of 4,591,276 molecules. It is becoming a crucial tool for exploring molecular space, where models learn from large datasets to produce novel molecular structures with similar properties. These structures are useful for virtual screening and training semi-supervised predictive models for downstream tasks.
  • ADMETlab 2.0 (Xiong et al., 2021). ADMETlab 2.0 is an advanced online platform dedicated to the accurate prediction of ADMET (Absorption, Distribution, Metabolism, Excretion, and Toxicity) properties. It hosts an extensive collection of bioactivity data, including 500,000 bioassay records and over 130 million bioactivity summary results, covering small molecules and RNAi reagents. This platform is pivotal in molecular modeling and drug discovery, offering a comprehensive framework for evaluating the ADMET profiles of chemicals.
4.3. Evaluation
To evaluate the capabilities of Mol-LLMs, we categorize them into four areas: property prediction, interaction prediction, reaction prediction, and molecule generation.
Property Prediction.
Molecular property prediction stands as a pivotal challenge in computational chemistry, where the goal is to accurately predict the properties of molecules, such as solubility, lipophilicity, affinity, absorption, distribution, metabolism, excretion, toxicity, and biological activity, based on their chemical structure. In this domain, most encoder-only models are typically pre-trained on extensive and unlabeled datasets and subsequently fine-tuned using specific data for property prediction. In the traditional domain of property prediction, BERT-based models such as ChemBERTa (Chithrananda et al., 2020), MAT (Maziarka et al., 2020), Mol-BERT  (Li and Jiang, 2021) and MG-BERT (Zhang et al., 2021) are widely recognized. However, in cases of small sample problems, SMILES Transformer (Honda et al., 2019) has demonstrated superior performance.
Interaction Prediction. Drug-drug interaction (DDI) prediction is a crucial field in drug discovery. Drug-drug interaction (DDI) prediction refers to the process of identifying and assessing the potential effects that may occur when two or more drugs are administered concurrently. This predictive analysis is essential for understanding how different pharmaceutical compounds might interact within the body. In the domain of DDI prediction, a representative model is X-Mol (Xue et al., 2020), which enhances the accuracy and reliability of DDI predictions, facilitating the development of safer drug combinations and reducing the risk of adverse reactions in patients.
Reaction Prediction. Chemical reaction prediction is a significant area of study in chemistry, and it generally falls into two main categories: forward reaction prediction and retrosynthetic analysis. The forward reaction prediction involves predicting the products of a chemical reaction based on the reactants. It’s about understanding how different chemicals react with each other under various conditions. Forward reaction prediction helps chemists anticipate the outcome of a chemical reaction, which is crucial for synthesizing new compounds, developing pharmaceuticals, and various industrial processes. In the domain of forward reaction prediction, Molecular Transformer (Schwaller et al., 2019) and GO-PRO (Mann and Venkatasubramanian, 2021) are the two classical models. The retrosynthetic analysis is the reverse of forward reaction prediction. It starts with a complex product and works backward to deduce the simpler starting materials (reactants) and the reactions needed to synthesize the product. This approach is particularly valuable in organic chemistry for designing pathways to synthesize complex organic molecules. It helps chemists identify the most efficient and practical way to produce a desired compound, often used in drug development and the synthesis of new materials. Retrosynthesis Transformer (Toniato et al., 2023) and SCROP (Zheng et al., 2019) are representative models in this domain. However, retrosynthesis prediction is challenging since there are massive possible synthetic routes available, and it is often difficult to navigate the direction of the retrosynthesis process. Therefore, there are GET (Mao et al., 2021) and Graph2SMILES (Tu and Coley, 2022) that incorporate critical topological connections between atoms, and Disconnection aware model (Thakkar et al., 2023) and RetroSynth-Diversity (Toniato et al., 2023) that integrate prompts.
Molecule Generation.
Molecule Generation is an emerging and critical area in computational chemistry and drug design, harnessing the power of advanced computational techniques to create novel molecular structures. This field is particularly vital in the development of new pharmaceuticals and materials, where the ability to generate molecules with desired properties can greatly accelerate the discovery process. Molecule generation encompasses two primary approaches: template-based design and de novo design. The former involves modifying known molecules or scaffolds. This approach starts with existing molecular frameworks, which are then altered to enhance desirable properties or reduce unwanted effects. The de novo design refers to the creation of novel molecular structures from scratch, without relying on pre-existing templates. This method is particularly valuable for discovering unique compounds with potential therapeutic or material applications. It challenges computational models to explore vast chemical spaces, requiring sophisticated algorithms capable of predicting viable and effective molecular structures. To address this task, there are lots of typical Mol-LLMs such as MolGPT (Bagal et al., 2021), SMILES GPT (Adilov, 2021) and Taiga (Mazuz et al., 2023).
Evaluation Metric To quantify the performance of Mol-LLMs in property prediction, interaction prediction, and reaction prediction tasks, which primarily involve classification and regression, one can employ the standard metrics in machine learning, such as accuracy, F1-score, and correlation coefficient. For molecular generation tasks, their performance is evaluated based on their ability to produce diverse, realistic, and relevant molecules. We present several popular metrics as follows.
  • Validity (Polykovskiy et al., 2020; Brown et al., 2019). This metric assesses the percentage of generated molecules that are chemically viable. Validity is crucial in ensuring that the output of generative models adheres to fundamental chemical laws, such as proper atomic valences and bond formations.
  • Uniqueness (Polykovskiy et al., 2020; Brown et al., 2019). It focuses on the uniqueness of the generated SMILES strings. It is calculated for the first $K$ valid molecules in the generated set, where $K$ is typically 1,000 or 10,000, ensuring the model’s capability to produce a broad range of distinct molecular structures. This metric is particularly vital in the context of high-dimensional chemical space, as it assesses the model’s ability to generate a vast array of unique molecules.
  • Novelty (Polykovskiy et al., 2020; Brown et al., 2019). Novelty is the fraction of the generated molecules that are not present in the training set. It measures the proportion of generated molecules that are unique compared to the training dataset. Low novelty indicates overfitting. This metric is significant in drug discovery, where the generation of novel compounds can lead to the development of new therapeutics.
[…]
    (1) $\text{IntDiv}_{p}(G)=1-\sqrt[p]{\left(\frac{1}{|G|^{2}}\sum_{m_{1},m_{2}\in G}T(m_{1},m_{2})^{p}\right)}.$
[…]
  • External diversity (Benhenda, 2017). The external diversity of a generative model is defined as the relative diversity between the training set and a sufficiently large generated sample. $A_{1}$ and $A_{2}$ are two sets of molecules, the relative diversity $E$ of $A_{1}$, $A_{2}$ is defined by:
    (2) $E\left(A_{1},A_{2}\right)=\frac{1}{\left|A_{1}\right|\times\left|A_{2}\right|}\sum_{(x,y)\in A_{1}\times A_{2}}T_{d}(x,y).$
    $T_{d}(x,y)$ is the Tanimoto distance between molecule $x$ from set $A_{1}$ and molecule $y$ from set $A_{2}$. The higher the external diversity, the more distinct the generated set is from the reference set, indicating a broader exploration of the chemical space by the generative model.
  • Fragment similarity (Frag) (Polykovskiy et al., 2020). This metric serves as a comparative tool for evaluating the distribution of BRICS fragments within generated and reference molecular sets. Note that BRICS is a method used to fragment molecules into smaller, more manageable pieces based on retrosynthetically interesting chemical substructures. The Frag metric is formulated by calculating the cosine similarity between the substructural elements of generated molecules and those within a reference set,
    (3) $Frag(G,R)=\frac{\sum\limits_{f\in F}[c_{f}(G)\cdot c_{f}(R)]}{\sqrt{\sum\limits_{f\in F}c^{2}_{f}(G)}\sqrt{\sum\limits_{f\in F}c^{2}_{f}(R)}},$
    where $c_{f}(A)$ represents the frequency of occurrence of a substructure $f$ within a set of molecules $A$, and $F$ denotes the collective set of fragments present in either the generated set $G$ or the reference set $R$.
  • Scaffold similarity (Scaff) (Polykovskiy et al., 2020). This metric is similar to the fragment similarity metric. However, instead of comparing the frequencies of fragments, it compares the frequencies of Bemis-Murcko scaffolds between the generated and reference sets of molecules. Note that Bemis-Murcko scaffolds include all of a molecule’s ring structures and linker fragments connecting rings.
  • Fréchet ChemNet Distance (FCD) (Polykovskiy et al., 2020; Brown et al., 2019). It measures the difference in distributions between a set of generated molecules $G$ and a reference set $R$ by comparing their mean vectors and covariance matrices of activations. These activations represent chemical and biological properties captured from the molecules. The formula for FCD is defined as follows:
    (4) $\text{FCD}(G,R)=\|\mu_{G}-\mu_{R}\|^{2}+\text{Tr}[\Sigma_{G}+\Sigma_{R}-2(\Sigma_{G}\Sigma_{R})^{1/2}].$
    where $\mu_{G}$ and $\mu_{R}$ are the mean vectors of activations for the molecules in sets $G$ and $R$ respectively, while $\Sigma_{G}$ and $\Sigma_{R}$ are the corresponding covariance matrices. A low FCD value indicates generated molecules closely resemble those in the reference set, both chemically and biologically, suggesting high-quality molecular generation.
4.4. Summary
In this section, we have meticulously examined the landscape of Sci-LLMs specifically adapted to chemical molecular languages like SMILES and SELFIES. Our exploration encompasses an in-depth analysis of their unique model architectures and capabilities, focusing on how these models interpret and process chemical languages. We also delve into the datasets that are pivotal for training and benchmarking these models, highlighting the criteria that are essential for evaluating their performance, particularly in the context of the increasingly popular field of molecular generation.
5. Protein Large Language Models
In the past years, large language models have become increasingly influential in protein research, offering novel insights and capabilities in understanding and manipulating proteins. In this section, we present a comprehensive review of LLMs for proteins (named Prot-LLMs), encompassing detailed discussions on their model architectures, utilized datasets, various capabilities, and corresponding evaluation criteria. The overview of this section is shown in Figure 9.
{forest}
for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw, rounded corners,align=left, minimum width=2.5em, inner xsep=4pt, inner ysep=1pt, , where level=1text width=5em,fill=blue!10, where level=2text width=5em,font=,fill=pink!30, where level=3font=,yshift=0.26pt,fill=yellow!20, [Prot-LLMs, fill=green!10 [Models
(Sec. 5.1),text width=4em [Encoder-only, text width= 5.5em [Sequence-only: ProteinBert (Brandes et al., 2022)   ProtTrans (Elnaggar et al., 2022)  ESM-1b (Rives et al., 2019)
ESM-2 (Lin et al., 2022)  PMLM  (He et al., 2021)   ProtFlash (Wang et al., 2023e)   ProteinNPT  (Notin et al., 2023b) ] [MSA-combined: ESM-MSA-1b (Rao et al., 2021)  ESM-1v (Meier et al., 2021)   RSA (Ma et al., 2023) ] [Add 3D structure: ESM-GearNet (Zhang et al., 2023d)  SaProt (Su et al., 2023)  LM-GVP (Wang et al., 2022a)  Mansoor et al. (Mansoor et al., 2021) ] [Integrating multi-level structure: PromptProtein (Wang et al., 2023d)  ] [Add extra knowledge: OntoProtein (Zhang et al., 2022)  KeAP (Zhou et al., 2023) ] ] [Decoder-only, text width= 5.5em [GPT-based: ProGen  (Madani et al., 2020)  ProGen2 (Nijkamp et al., 2023)   RITA (Hesslow et al., 2022)  ProtGPT2 (Ferruz et al., 2022)
LM-D Design (Zheng et al., 2023a)  PoET (au2 and Bepler, 2023)  IgLM (Shuai et al., 2021)  ZymCTRL (Munsamy et al., 2022) ] ] [Encoder-decoder, text width=5.5em [T5-based: ProstT5  (Heinzinger et al., 2023)  pAbT5  (Chu and Wei, 2023) ] [Transformer-based: SS-pLM (Serrano et al., 2023) ] [GLM-based: xTrimoPGLM (Chen et al., 2023a) ] [MSA-combined: MSA2Prot  (Ram and Bepler, 2022)  MSA-Augmenter  (Zhang et al., 2023)  EvoOpt (Yamaguchi and Saito, 2022)  Sgarbossa et al. (Sgarbossa et al., 2023) ] [Others: Fold2Seq (Cao et al., 2021)  Lee et al. (Lee et al., 2022) ] ] ] [Datasets
(Sec. 5.2),text width=4em [Pre-training, text width=4em [UniRef100  UniRef90  UniRef50 (Suzek et al., 2007, 2015)  Swiss-Prot (Boutet et al., 2016)  TrEMBL (Moller et al., 1999)
UniParc (Consortium, 2023)  Pfam (Finn et al., 2006)  BFD (Steinegger and Söding, 2018; Steinegger et al., 2019; Jumper et al., 2021)  PDB (wwPDB consortium, 2018)  AlphaFoldDB (Jumper et al., 2021; Varadi et al., 2021) ] ] [Benchmarks, text width=4em [CASP (Kryshtafovych et al., 2021)   EC (McDonald et al., 2008)  GO (Ashburner et al., 2000)  CATH (Orengo et al., 1997)  HIPPIE (Schaefer et al., 2012)  SCOP (Lo Conte et al., 2000)
ProteinGym (Notin et al., 2023a)  FLIP (Dallago et al., 2021)  PEER (Xu et al., 2022)  TAPE (Rao et al., [n. d.]) ] ] ] [Evaluation
(Sec. 5.3),text width=4em [Structure Prediction, text width=7em [Secondary  tertiary and quaternary structure predictions] ] [Function Prediction, text width=7em [ Protein classification; Protein-protein interaction; Protein-ligand interaction;
Subcellular localization; Binary localization; Remote homology detection;
Fluorescence; Stability; $\beta$-Lactamase activity; Solubility; Mutation effect prediction ] ] [Sequence Generation, text width=7em [ De novo protein design; Protein sequence optimization] ] [Evaluation metric, text width=7em [ Structure prediction tasks: RMSD GDT TM-Score LDDT etc. ] [ Function prediction tasks: Accuracy Recall F1-score etc. ] [ Sequence generation tasks: PPL Novelty FPD Diversity Foldability Recovery ] ] ] [Summary
(Sec. 5.4),text width=4em ] ]
Figure 9. Chapter overview of Prot-LLMs.
5.1. Models
In this section, we also classify Prot-LLMs into three main types based on their specific architectures: encoder-only, decoder-only, and encoder-decoder models. These distinct architectures are well-suited for diverse protein research applications. For example, the encoder-only models primarily serve protein function or property prediction purposes, while the decoder-only models are predominantly employed for protein generation tasks. Table 5 provides a summary of Prot-LLMs.
5.1.1. Encoder-only Models
Most of the encoder-only Prot-LLMs are built upon the encoder of Transformer, which enables the encoding of protein sequences or structures into fixed-length vector representations. These representations facilitate the identification of patterns and features within proteins, thereby enhancing subsequent analysis and prediction tasks. Self-supervised learning is employed to acquire protein representation, such as masked language modeling (MLM) tasks that reconstruct corrupted tokens based on the surrounding sequence. Prominent pre-trained protein sequence encoders include ESM-1b (Rives et al., 2019), ESM-1v (Meier et al., 2021), and ESM-2 (Lin et al., 2022), ProteinBert (Brandes et al., 2022), ProtTrans (Elnaggar et al., 2022). The ESM series (Rives et al., 2019; Meier et al., 2021; Lin et al., 2022) mainly employs a Transformer’s encoder (i.e., BERT (Devlin et al., 2018) and RoBERTa (Liu et al., 2019)) to predict protein structure and function, which utilizes extensive sequence information from protein databases without relying on manual annotations of the sequences. ProteinBert (Brandes et al., 2022) enhances the classical BERT architecture by incorporating a novel pretraining task for predicting protein functionality. It separates local (character-level) and global (sequence-level) representations, enabling multitask processing in a principled way. ProtTrans (Elnaggar et al., 2022) trained several auto-encoder models (includes BERT (Devlin et al., 2018), Albert (Lan et al., 2019), Electra (Clark et al., 2020)) on a vast of sequence data. Several works have attempted to improve the encoder architecture and the training method. PMLM (He et al., 2021) introduced Pairwise Masked Language Model (PMLM), to directly pretrain the encoder with a focus on capturing co-evolutionary information reflected in residue co-variation within the sequence. This approach considers dependencies between masked tokens. ProtFlash (Wang et al., 2023e) suggested a mixed-chunk attention mechanism that combines multiple positional encodings and a hybrid block attention mechanism incorporating both local and global attention, effectively reducing model complexity. ProteinNPT (Notin et al., 2023b) developed a non-parametric Transformer specifically designed for protein sequences, making it particularly suitable for scenarios involving sparse labels and multi-task learning.
The Multiple Sequence Alignment (MSA) is a computational method that can reveal common features and variation patterns among sequences. By aligning multiple sequences, MSA can unveil shared evolutionary relationships between them and facilitate the identification of functional regions and structural domains. Therefore, MSA has been widely used in protein models. For example, ESM-MSA-1b (MSA Transformer) (Rao et al., 2021) extends the self-attention mechanism to the MSA setting, which interleaves self-attention across rows and columns to capture dependencies between amino acids and between sequences. Notably, MSA Transformer has become a crucial component of AlphaFold2 (Fang et al., 2023b) and AlphaMissence (Cheng et al., 2023). However, the utilization of MSA incurs significant computational overheads, and cannot handle orphan proteins. An important alternative method, called Retrieved Sequence Augmentation (RSA) (Ma et al., 2023), eliminates the need for additional alignment or pre-processing steps. RSA effectively links query protein sequences to a set of structurally or functionally similar sequences in the database and aggregates these sequences for subsequent function prediction tasks.
Given that the function of a protein is intricately linked to its structure, another reasonable approach for protein representation would involve the integration of its 3D conformation. There are some recent efforts in this direction. ESM-GearNet (Zhang et al., 2023d) combines ESM-1b and GearNet (Xie et al., 2022) (the state-of-the-art protein structure encoder) to enhance model performance. SaProt (Su et al., 2023) introduces a structure-aware vocabulary that integrates residue tokens with structure tokens. LM-GVP (Wang et al., 2022a) compose Transformer blocks and a graph network derived from the protein 3D structure. Mansoor et al. (Mansoor et al., 2021) encoded protein sequence and structure through joint training in a semi-supervised manner. PromptProtein (Wang et al., 2023d) utilizes prompt-guided multi-task pretraining to fuse different levels of protein structure. It is worth noting that numerous 3D structure encoders (like GearNet (Xie et al., 2022) and GBPNet (Aykent and Xia, 2022)) have been proposed to model protein structural information; however, they are beyond the scope of this survey which primarily focuses on LLMs.
Additionally, some approaches propose to introduce extra knowledge such as Gene Ontology (GO) to enhance the protein representation. OntoProtein (Zhang et al., 2022) considered GO as a factual knowledge graph, which was used to enhance protein representations. KeAP (Zhou et al., 2023) proposed a knowledge-injected auto-encoder that performs token-level knowledge graph exploration for protein representation learning.
Table 5. Summary of Prot-LLMs
Model Time #Parameters Base model Pretraining Dataset Capability Open-source
Encoder-only ESM-1b  (Rives et al., 2019) 2020.02 650M RoBERTa UniRef50 Secondary struct. pred., Contact pred., etc.
ESM-MSA-1b  (Rao et al., 2021) 2021.02 100M ESM-1b UniRef50 Secondary struct. pred., Contact pred., etc.
ESM-1v  (Meier et al., 2021) 2021.02 650M ESM-1b UniRef90 Mutation effect pred.
ProtTrans  (Elnaggar et al., 2022) 2021.07 - BERT, Albert, Electra UniRef, BFD Secondary struct. pred., Func. pred., etc
PMLM (He et al., 2021) 2021.07 87M - 731M Trans. enc. Uniref50/Pfam Contact pred. $\times$
Mansoor et al.  (Mansoor et al., 2021) 2021.09 100M ESM-1b - Mutation effect pred. $\times$
ProteinBERT  (Brandes et al., 2022) 2022.02 16M BERT UniRef90 Func. pred.
LM-GVP  (Wang et al., 2022a) 2022.04 - Trans. enc - Func. pred.
RSA  (Ma et al., 2023) 2022.05 - ESM-1b - Func. pred.
OntoProtein  (Zhang et al., 2022) 2022.06 - BERT ProteinKG25 Func. pred.
ESM-2  (Lin et al., 2022) 2022.07 8M - 15B RoBERTa UniRef50 Func. pred., Struct. pred.
PromptProtein  (Wang et al., 2023d) 2023.02 650M RoBERTa UniRef50, PDB Func. pred.
KeAP  (Zhou et al., 2023) 2023.02 - RoBERTa ProteinKG25 Func. pred.
ProtFlash  (Wang et al., 2023e) 2023.10 79M/174M Trans. enc UniRef50 Func. pred.
ESM-GearNet  (Zhang et al., 2023d) 2023.10 - ESM-1b, GearNet - Func. pred.
SaProt  (Su et al., 2023) 2023.10 650M BERT - Mutation effect pred.
ProteinNPT  (Notin et al., 2023b) 2023.12 - Trans. enc. - Fitness pred., Redesign $\times$
Decoder-only ProGen  (Madani et al., 2020) 2020.03 1.2B GPT Uniparc SWISS-Prot Functional prot. gen.
ProtGPT2  (Ferruz et al., 2022) 2021.01 738M GPT Uniref50 De novo protein design and engineering
ZymCTRL  (Munsamy et al., 2022) 2022.01 738M GPT BRENDA Functional enzymes gen.
RITA  (Hesslow et al., 2022) 2022.05 1.2B GPT UniRef100 Functional prot. gen. $\times$
IgLM  (Shuai et al., 2021) 2022.12 13M GPT - Antibody design
LM-D  (Zheng et al., 2023a) 2023.02 664M GPT - Enzyme design $\times$
ProGen2  (Nijkamp et al., 2023) 2023.10 151M - 6.4B GPT Uniref90, BFD30, PDB Functional prot. gen.
PoET  (au2 and Bepler, 2023) 2023.11 201M GPT - Prot. family. gen. $\times$
Encoder-Decoder Fold2Seq  (Cao et al., 2021) 2021.01 - Transformer - Prot. design
MSA2Prot  (Ram and Bepler, 2022) 2022.04 - Transformer - Prot. gen., Variant func. pred. $\times$
Sgarbossa et al. (Sgarbossa et al., 2023) 2023.02 - MSA Transformer - Prot. gen.
Lee et al. (Lee et al., 2022) 2023.02 150M Transformer - Prot. design $\times$
MSA-Augmenter  (Zhang et al., 2023) 2023.06 260M Transformer Uniref50 MSA gen.
ProstT5  (Heinzinger et al., 2023) 2023.07 3B T5 PDB Seq.-struct. translation
xTrimoPGLM  (Chen et al., 2023a) 2023.07 100B GLM Uniref90, ColdFoldDB Prot. gen., Func. pred. $\times$
SS-pLM (Serrano et al., 2023) 2023.08 14.8M Transformer Uniref50 Prot. gen. $\times$
pAbT5  (Chu and Wei, 2023) 2023.10 - T5 - Prot. design $\times$
5.1.2. Decoder-only Models
Decoder-based Prot-LLMs play a predominant role in the generation of novel proteins, serving as a crucial tool in protein engineering and drug design. GPT (Radford et al., 2019b), a well-established decoder-only architecture, has gained extensive utilization in Prot-LLMs. A representative model is ProGen  (Madani et al., 2020), which utilizes GPT for controllable protein generation. It has been trained on a dataset comprising 280M protein sequences, accompanied by conditioning tags that encode diverse annotations encompassing taxonomic, functional, and locational information. ProGen2 (Nijkamp et al., 2023) extends the model to 6.4B parameters and is trained on diverse sequence datasets extracted from over one billion proteins from genomic, metagenomic, and immune repertoire databases. ProtGPT2 (Ferruz et al., 2022) is another GPT-based autoregressive model that effectively generates protein sequences exhibiting amino acid compositions and disorder propensities comparable to those observed in natural proteins. RITA (Hesslow et al., 2022) presents a suite of autoregressive generative models, which is the first systematic investigation of how capabilities evolve with model size for these protein models.
The decoder-based protein language models have been widely applied in specific protein designs. Timothy et al. introduced the protein evolutionary Transformer (PoET), an autoregressive generative model of the distribution over protein families, which can generate sets of related proteins. Zheng et al. proposed the LM-Design (Zheng et al., 2023a) that uses language models for structure-based protein design (i.e., inverse folding). ZymCTRL (Munsamy et al., 2022) trained a conditional language model on the BRENDA enzyme database (Chang et al., 2021), aimed at designing customized artificial enzymes by generating specific enzyme classes based on user prompts. IgLM (Shuai et al., 2021) employed autoregressive sequence generation for antibody design.
5.1.3. Encoder-decoder Models
The encoder-decoder architecture is commonly used for sequence-to-sequence tasks. In the field of protein research, ProstT5  (Heinzinger et al., 2023), a protein language model based on the T5 architecture (Raffel et al., 2020), is an exemplary encoder-decoder model that facilitates translation between protein sequences and structures. The conversion of protein structure from 3D to 1D is achieved through the utilization of 3Di-tokens introduced by Foldseek (van Kempen et al., 2022). pAbT5 (Chu and Wei, 2023) is another T5-based model that takes into consideration the constraints imposed by protein-protein interactions on generation and design. It can generate complementary heavy or light chains from their pairing partners. xTrimoPGLM (Chen et al., 2023a), based on General Language Model (GLM) architecture (Du et al., 2021), is a unified protein language model for understanding and generating protein sequences. It has an unprecedented scale of 100B parameters and consumes 1 trillion training tokens. Considering that the training of protein LLMs involves millions to billions of sequences and billions of parameters, Serrano et al. (Serrano et al., 2023) introduced the Small-Scale protein Language Model (SS-pLM), which significantly reduces the computational burden, democratizing the use of foundational models in protein generation.
Building upon the Transformer, several models have incorporated MSA modules. MSA2Prot  (Ram and Bepler, 2022) is an MSA-to-protein Transformer, developed axial and cross attentions for encoder and decoder to model sequence probabilities autoregressively. MSA-Augmenter  (Zhang et al., 2023) utilizes protein-specific attention mechanisms and large-scale MSAs to generate novel protein sequences that are not present in existing databases. Sgarbossa et al. (Sgarbossa et al., 2023) proposed an iterative method that directly utilizes the masked language modeling to generate sequences within the MSA Transformer framework.
In the field of structure-based protein design, Cao et al. (Cao et al., 2021) proposed a Fold2Seq framework, a generative approach based on the Transformer architecture, which aims at designing protein sequences conditioned on specific target folds. To capture intricate sequence-structure relationships, Fold2Seq jointly learns embeddings for both sequences and folding structures, which are derived from the density of secondary structure elements in 3D voxels. Additionally, some approaches utilize reinforcement learning to improve the generation quality. Lee et al. (Lee et al., 2022) efficiently navigates the latent representation space rather than the protein sequence space by formulating the sequence design task as a Markov decision process and employing model-based reinforcement learning. This approach enables them to generate proteins with enhanced functionality and cellular fitness.
Table 6. Summary of datasets for Prot-LLMs
Dataset Time Scale Keywords
Pretraining UniRef100 (Suzek et al., 2007, 2015) 2023.11 314M Complete collection of protein sequences from UniProtKB
[…]
UniProtKB/Swiss-Prot (Boutet et al., 2016) 2023.11 570K High-quality, manually curated protein sequence database
UniProtKB/TrEMBL (Moller et al., 1999) 251M Computationally annotated protein sequence database
UniParc (Consortium, 2023) 2023.11 632M Comprehensive and non-redundant protein sequence database
Pfam (Finn et al., 2006) 2023.09 47M Protein family database
BFD (Steinegger and Söding, 2018; Steinegger et al., 2019; Jumper et al., 2021) 2021.07 2.5B Protein sequences from multiple databases and resources
PDB (wwPDB consortium, 2018) 2023.12 214K Experimentally determined accurate protein structures
AlphaFoldDB (Jumper et al., 2021; Varadi et al., 2021) 2021.11 200M Protein structures predicted by AlphaFold
Benchmark CASP (Kryshtafovych et al., 2021) 2022.01 - Structure prediction competition
EC (McDonald et al., 2008) 2023.11 2.6 M Enzymes classification database
GO (Ashburner et al., 2000) 2023.11 1.5M Gene Ontology knowledgebase
CATH (Orengo et al., 1997) 2023.02 151M Classification of protein structures
HIPPIE (Schaefer et al., 2012) 2022.04 39K Protein-protein interaction networks
SCOP (Lo Conte et al., 2000) 2023.01 914K Protein structure classification
ProteinGym (Notin et al., 2023a) 2022.12 $\sim 300$K Predict the effects of protein mutations
FLIP (Dallago et al., 2021) 2022.01 $\sim 320$K Fitness landscape prediction (AAV, Thermostability, GB1)
PEER (Xu et al., 2022) 2022.11 $\sim 390$K Protein function, Localization, Structure prediction, Protein-protein interaction, Protein-ligand interaction
TAPE (Rao et al., [n. d.]) 2021.09 $\sim 120$K Remote homology detection, Secondary structure, Contact, Fluorescence, Stability prediction
5.2. Datasets
Protein datasets can be classified into two categories based on the availability of annotations: pre-training datasets and benchmarks. The former, which lacks labels, is commonly employed for self-supervised pre-training, whereas the latter, with labeled data, is utilized for supervised fine-tuning or model evaluation. The following present popular pre-training datasets and benchmarks for Prot-LLMs, as summarized in Table 6.
Pre-training Datasets:
  • UniProtKB (Swiss-Prot and TrEMBL) (Boutet et al., 2016; Moller et al., 1999). It is a central hub for the collection of functional information on proteins, with extensive and accurately annotated protein sequence data. It comprises Swiss-Prot and TrEMBL, where the former is known for its high level of annotations, accuracy, and reliability, because they are manually annotated and reviewed by experts. TrEMBL contains protein sequences that are automatically annotated and have not yet been reviewed by human curators.
  • UniRef100, UniRef90, UniRef50 (Suzek et al., 2007, 2015). UniRef (UniProt Reference Clusters) is a system of databases that provide clustered sets of protein sequences from UniProtKB and selected UniProt Archive records. The primary purpose of UniRef is to provide a comprehensive and non-redundant dataset that improves the efficiency of sequence similarity searches. The databases are organized into three main clusters: UniRef100, UniRef90, and UniRef50, where the latter two are created by clustering UniRef100 sequences that have at least 90% and 50% sequence identity to each other, respectively.
  • UniParc (Consortium, 2023). UniParc (UniProt Archive) is an important component of the UniProt database family, serving as a resource for storing non-redundant protein sequences. Unlike other members of the UniProt database family, UniParc does not provide protein annotation and functional information; it focuses solely on storing protein sequences themselves.
  • Pfam (Finn et al., 2006). Pfam is a widely used protein family database that is used for the classification and annotation of protein sequences. It plays a crucial role in understanding protein function, structure, and evolution. It is built based on the shared conserved domains and functional modules found in protein sequences, providing researchers with classification, annotation, and functional prediction of protein sequences.
  • BFD (Steinegger and Söding, 2018; Steinegger et al., 2019; Jumper et al., 2021). The Big Fantastic Database (BFD) is a comprehensive collection of protein sequences gathered from various sources. It includes a vast array of sequence data (2.5B), which is crucial for training machine learning models in tasks like protein structure prediction.
  • PDB (wwPDB consortium, 2018).The Protein Data Bank (PDB) is a global protein structure database used for storing and sharing resolved 3D protein structure data. It archives detailed information about the 3D structures of proteins, nucleic acids, and complex assemblies. These structures are determined using experimental methods like X-ray crystallography, NMR spectroscopy, and, more recently, cryo-electron microscopy.
  • AlphaFoldDB  (Jumper et al., 2021; Varadi et al., 2021) AlphaFold Database comprises protein structure predictions generated by DeepMind’s AlphaFold system, offering accurate predictions of protein 3D structures. While these structures are not experimentally determined, they offer high-confidence models that can be incredibly useful, especially for proteins whose structures have not yet been solved experimentally.
Benchmarks:
  • CASP (Kryshtafovych et al., 2021). The Critical Assessment of Structure Prediction (CASP) is an influential initiative in the field of protein structure prediction. The goal of CASP is to advance the methods used in predicting protein structures from their amino acid sequences. CASP is essentially a series of competitions, which provides a unified benchmark for evaluating the performance of different methods in protein structure prediction.
  • EC (McDonald et al., 2008). The Enzyme Commission (EC) dataset is a standardized dataset used to describe the classification and functions of enzymes. Enzymes are protein molecules that are widely present in living organisms and play a crucial role in catalyzing chemical reactions. The EC dataset contains a large number of enzyme records, each with a unique EC number and corresponding description. These descriptions include the enzyme’s name, reaction type, catalytic substrates, and products.
  • GO (Ashburner et al., 2000). The Gene Ontology (GO) dataset is a widely used gene annotation and functional classification system for describing the biological functions of genes and proteins. It provides a standardized framework for representing gene and gene product attributes across species and databases, helping researchers understand their functions in terms of cellular processes, molecular functions, and cellular components.
  • CATH (Orengo et al., 1997). The acronym CATH stands for the four main levels at which it classifies protein domains: Class, Architecture, Topology, and Homologous superfamily. CATH is widely used for protein structure classification and annotation. It offers a comprehensive framework for systematically classifying protein structures while providing crucial insights into their function and evolutionary aspects.
  • SCOP (Lo Conte et al., 2000). SCOP (Structural Classification of Proteins) is a database used for protein structure classification. It helps researchers understand the structure, function, and evolutionary relationships of proteins by assigning protein structures to different levels and categories. The SCOP database employs a hierarchical classification system consisting of four levels: Class, Fold, Superfamily, and Family. Each level has specific definitions and criteria to group protein structures into their respective categories.
  • ProteinGym (Notin et al., 2022). The ProteinGym platform comprises a comprehensive collection of Deep Mutational Scanning (DMS) assays meticulously curated to evaluate mutation effect predictors in determining the fitness of mutated proteins. It encompasses two distinct benchmarks: a substitution benchmark encompassing experimental characterization of 1.5 million missense variants across 87 DMS assays, and an indel benchmark comprising analysis of 300,000 mutants across 7 DMS assays.
  • FLIP (Dallago et al., 2021). FLIP (Functional Learning for Protein Engineering) is a benchmark designed to encourage rapid scoring of representation learning for protein engineering. Currently, FLIP covers experimental data on adenovirus stability for gene therapy, stability of protein domain B1, immunoglobulin binding, and thermal stability of multiple protein families.
  • PEER (Xu et al., 2022). The PEER benchmark is a comprehensive and multi-task benchmark for protein sequence understanding, encompassing 17 tasks across five categories: protein function, localization, structure predictions, protein-protein interaction prediction and protein-ligand interaction prediction. This benchmark evaluates diverse sequence-based methodologies for each task, including conventional feature engineering approaches, various sequence encoding methods, as well as large-scale pre-trained protein language models under both single-task learning and multi-task learning settings.
  • TAPE (Rao et al., [n. d.]). The Tasks Assessing Protein Embeddings (TAPE) dataset is a collection of five biologically relevant semi-supervised learning tasks spanning different domains of protein biology, including secondary structure prediction, contact prediction, remote homology prediction, stability prediction, and fluorescence prediction. TAPE divides the tasks into specific training, validation, and testing sets to ensure the evaluation of each task’s biological relevance and generalizability to real-life scenarios. TAPE serves as a benchmark for a range of semi-supervised protein representation learning methods, encompassing recent advances as well as established sequence-based learning techniques.
5.3. Evaluation
In the realm of protein computational models, Prot-LLMs are primarily evaluated based on their capabilities in three critical areas: protein function prediction, protein sequence generation, and protein structure prediction. Protein Structure Prediction. It is a crucial area within the fields of bioinformatics and computational biology. It involves the prediction of the three-dimensional structure of proteins given an input sequence, which includes determining the atomic coordinates and the topological relationships between atoms. Predicting protein structures is of significant importance for understanding protein function, drug design, and biomedical research. This task includes tertiary and quaternary structure predictions. The former is critical for understanding the functional mechanisms of a protein, and the latter plays a vital role in understanding the functionality of multi-subunit complexes like hemoglobin. The models for protein structure prediction often rely on encoder-based Prot-LLMs (i.e., ESM series (Rives et al., 2019; Meier et al., 2021; Lin et al., 2022)) to extract the sequence information, followed by structure prediction modules.
Protein Function Prediction. This task aims to determine the biological function of proteins, including how they operate within an organism and their interactions with other biomolecules. The prediction of protein function encompasses a multitude of subtasks, enumerated below.
  • Protein Classification. This refers to the process of categorizing and grouping proteins based on their structure, function, sequence similarity, or other features.
  • Protein-Protein Interaction Prediction. This refers to the focus on identifying and predicting interactions between different proteins. Proteins often interact with each other to perform various biological functions such as signal transduction, enzyme catalysis, and assembly of protein complexes.
  • Protein-Ligand Interaction Prediction. This refers to the process of predicting the binding interactions between a protein and a small molecule ligand.
  • Subcellular Localization Prediction. This refers to the prediction of the subcellular location of a protein within a cell. The subcellular localization of a protein is crucial for its function and interactions.
  • Binary Localization Prediction. This refers to the process of predicting the subcellular localization of a protein into one of two possible categories.
  • Remote Homology Detection. This refers to the process of identifying distant relationships between protein sequences. In the protein world, homology refers to the shared ancestry or evolutionary relationship between two or more proteins.
  • Fluorescence Landscape Prediction. This refers to the prediction of protein sequences to evaluate their fluorescence spectral characteristics under different environmental conditions.
  • Stability Prediction. This refers to the process of predicting the stability of a protein to assess its stability under specific conditions. Protein stability represents its ability to maintain the correct folding state within a biological organism or specific environment.
  • $\beta$-Lactamase Activity Prediction. This refers to the process of predicting the activity of a protein as a $\beta$-lactamase enzyme, which is capable of hydrolyzing $\beta$-lactam bonds and is commonly involved in antibiotic degradation and the development of antibiotic resistance.
  • Solubility Prediction. This refers to the process of predicting the solubility of a compound in a solution. Solubility represents the maximum amount of a compound that can dissolve in a solvent at a specific temperature and pressure.
  • Mutation Effect Prediction. It focuses on understanding how genetic mutations can affect protein function, a critical aspect due to their potential to induce diseases or modify biological processes. Prot-LLMs aim to identify potential mutation sites and predict the functional consequences of these mutations.
Protein Sequence Generation. It refers to the computational process of creating novel amino acid sequences that could potentially form functional proteins, which can be used for various applications such as drug design, enzyme engineering, and fundamental biological research. The task of protein sequence generation can be roughly classified into two categories:
  • De Novo Protein Design. It involves creating entirely new protein sequences that are not based on existing proteins. This process relies heavily on computational methods to design proteins that can perform desired functions, such as binding to a particular target. The autoregressive generative models (e.g., ProGen series) are often employed for the tasks of protein sequence generation.
  • Protein Sequence Optimization. It focuses on modifying existing protein sequences to enhance or alter their functions, employing techniques such as directed evolution. Directed evolution is particularly useful for improving the properties of existing proteins, such as increasing their stability, altering their substrate specificity, or enhancing their catalytic efficiency.
Evaluation metric. The evaluation of protein structure prediction models typically includes Root Mean Square Deviation (RMSD), Global Distance Test (GDT) (Zhang and Skolnick, 2007), Template Modeling Score (TM-Score) (Zhang and Skolnick, 2007), and Local Distance Difference Test (LDDT) (Mariani et al., 2013). For more details on evaluating the predicted protein structure, readers can refer to (Kryshtafovych et al., 2021). For protein function prediction, the standard metrics in machine learning (e.g., accuracy for classification tasks and correlation coefficient for regression tasks) can be employed to quantitatively assess the performance of these Prot-LLMs. To evaluate the performance of protein sequence generation, one can use the following metrics: the model perplexity, novelty, similarity, and diversity of generated protein sequences, as well as the condition consistency if the generation is conditioned on certain inputs (e.g., protein backbone). We list them as follows.
  • Perplexity (PPL). In information theory, perplexity is a measure of uncertainty in the value of a sample from a discrete probability distribution. The larger the perplexity, the less likely it is that a language model can generate the protein sequence. The perplexity of a heldout test set reflects the capability of model to capture the distribution of natural sequences. The formula for PPL is defined as follows:
    (5) $\text{PPL}(p)=2^{H(p)}=2^{-\sum_{x}p(x)\log_{2}p(x)}=\Pi_{x}p(x)^{-p(x)},$
    where $H(p)$ is the entropy of the distribution $p$, and $x$ ranges over the protein sequence.
  • Novelty. Novelty refers to the uniqueness or novelty of a protein sequence in comparison to known proteins, which can be reflected by the highest identity between each generated sequence and natural proteins. The calculation involves aligning the amino acid sequences and counting the number of positions in which the amino acids are identical, defined as
    (6) $\text{Identity}(\%)=(\frac{N_{\text{ident}}}{N_{\text{total}}})\times 100,$
    where $N_{\text{ident}}$ represents the number of identical positions and $N_{\text{total}}$ represents the total number of aligned positions. A protein with a low percentage identity within all natural proteins implies a high degree of novelty.
  • Fréchet Protein Distance (FPD) (Jiang et al., 2008). This metric quantifies the similarity between a set of generated proteins ($G$) and a reference set ($R$). Previous research has demonstrated that protein language model embeddings inherently encapsulate information related to both structure and function, suggesting that the FPD metric provides a comprehensive measure of coverage across the distribution of structural and functional properties. The FPD is mathematically defined as follows:
    (7) $\text{FPD}=||\mu_{G}-\mu_{R}||^{2}+\text{Tr}(\Sigma_{G}+\Sigma_{R}-2\sqrt{\Sigma_{G}\Sigma_{R}}),$
    where, given the embedding space feature vectors for the reference and generated distributions, $\mu$ represents the feature-wise mean for each set of sequences, $\Sigma$ denotes the representative covariance matrix, and Tr refers to the trace linear algebra operation, defined as the sum of elements along the main diagonal of a square matrix.
  • Diversity. Sequence diversity assessment involves analyzing the variety of the generated proteins. Tools like BLAST can be used to compare these sequences against known protein databases to identify novel sequences. Metrics such as sequence similarity, percentage of unique sequences, and alignment scores provide a quantitative measure of diversity.
  • Foldability. This metric measures the foldability of a generated protein sequence by assessing the average per-residue confidence score, denoted as pLDDT, across the entire protein sequence. pLDDT serves as an indicator of the structural prediction model’s confidence in its predictions for individual residues. Lower pLDDT scores are frequently associated with intrinsically disordered regions (IDRs) within proteins, which lack a well-defined three-dimensional structure and thus exhibit inherent unfoldability.
  • Recovery. This metric refers to the success or accuracy in predicting the correct amino acid sequence that corresponds to a given 3D structure (i.e., Structure-based conditional generation). A high recovery rate indicates that the designed sequences are likely to fold the desired structures.
5.4. Summary
In this section, we have thoroughly examined the landscape of Sci-LLMs specifically designed for protein languages, namely Prot-LLMs. Our investigation encompasses a comprehensive analysis of their model architectures and capabilities, with a focus on how these models interpret and process protein languages. Additionally, we have curated an extensive collection of datasets, vital for training and benchmarking Prot-LLMs. These datasets are instrumental in providing a rich source of protein-related information, encompassing various aspects like sequence homology, structural configurations, and functional annotations. Finally, we present the evaluation method of Prot-LLMs, which encompasses several downstream tasks that these models are expected to perform, such as functional classification, interaction prediction, and protein sequence design. These tasks are crucial in assessing the models’ practical applicability in real-world biological and pharmaceutical scenarios.
6. Genomic Large Language Models
In the field of computational biology, genomic data exhibits resemblances to sequence-based information observed in natural language, enabling the application of large language models for analyzing genomic sequences. In this section, we provide a review of LLMs tailored for genomic language (Gene-LLMs), including insights into their model architectures, datasets, and evaluation. The overview of this section is shown in Figure 10.
{forest}
for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw, rounded corners,align=left, minimum width=2.5em, inner xsep=4pt, inner ysep=1pt, , where level=1text width=5em,fill=blue!10, where level=2text width=5em,font=,fill=pink!30, where level=3font=,yshift=0.26pt,fill=yellow!20, [Gene-LLMs, fill=green!10 [Models
(Sec. 6.1),text width=4em [Encoder-only, text width=4.5em [Conv series: GPN (Benegas et al., 2023b)   GPN-MSA (Benegas et al., 2023a)   LOGO (Yang et al., 2021)   ] [BERT variant:
iEnhancerBERT (Luo et al., 2022a)   iEnhancer-ELM (Li et al., 2023e)   miProBERT (Wang et al., 2023a)   TFBert (Luo et al., 2023a)   Enformer (Avsec et al., 2021)
BioSeq-BLM (Li et al., 2021)   DNABERT (Ji et al., 2021)   DNABERT-2 (Zhou et al., 2023b)  iDNA-ABF (Jin et al., 2022)   EpiGePT (Gao et al., 2023a)
Nucleotide-Transformer (Dalla-Torre et al., 2023)   SA DNA-LM (Gankin et al., 2023)   GENA-LM (Fishman et al., 2023)   PLPMpro (Li et al., 2023b)   Uni-RNA (Wang et al., 2023b) ] [Other series: MoDNA (An et al., 2022)   ] ] [Decoder-only, text width=4.5em [LongConv series: HyenaDNA (Nguyen et al., 2023)   ] [GPT series: DNAGPT (Zhang et al., 2023e)   ] [Diffusion+Transformer series: GenSLMs (Zvyagin et al., 2022)   ] ] [Encoder-Decoder, text width=6em [ ENBED (Malusare et al., 2023)   ] ] ] [Datasets
(Sec. 6.2),text width=4em [Pre-training, text width=4em [NCBI-Genome  UCSC-Genome  (Karolchik et al., 2003)  GRCh38  (Guo et al., 2017)   EPDnew (Dreos et al., 2013)   TAIR (Lamesch et al., 2012)
Mammalian-Gene (Strausberg et al., 1999)   ENCODE (coordination, data analysis coordination)   DeepSEA (Zhou and Troyanskaya, 2015)   VGDB (Hiscock and Upton, 2000)   Ensembl (Lamesch et al., 2012)   690 ChIP-seq (Zeng et al., 2016)
1000 Genomes Project (Consortium et al., 2015)   Panglao (Franzén et al., 2019)   ExPecto (Rusk, 2018)   BV-BRC (Olson et al., 2023)   RNAcmap (Rusk, 2018)   ] ] [Benchmark, text width=4em [CAGI5 (Katsonis and Lichtarge, 2019)   Protein-RNA (Horlacher et al., 2023)   NT-Bench (Dalla-Torre et al., 2023)   ] ] ] [Evaluation
(Sec. 6.3),text width=4em [Function Prediction, text width=7em [Promoter prediction] [Enhancer prediction] [Binding site prediction] ] [Structure Prediction, text width=7em [Chromatin profile prediction] [DNA/RNA-protein interaction prediction] ] [Sequence Generation, text width=7em ] [Sequence Variation and Evolution Analysis, text width=14em ] [Evaluation Metric, text width=7em [F1-score  Sensitivity  Specificity  MCC  AUROC] ] ] [Summary
(Sec. 6.4),text width=4em ] ]
Figure 10. Chapter overview of Gene-LLMs.
6.1. Models
Genomic LLMs, based on the Transformer architecture, proficiently model DNA or RNA sequence information by capturing long-range dependencies to accomplish prediction and generation tasks. By employing self-supervised learning techniques on the genomic sequences, Gene-LLMs gradually acquire an understanding of the genome. Once fine-tuned, or through contextual learning, they can be extensively utilized for downstream tasks, thereby improving accuracy and reducing the need for manual intervention. Here we categorize Gene-LLMs into three distinct types based on their specific architectures: encoder-only, decoder-only, and encoder-decoder models, as shown in Table 7.
Table 7. Summary of Gene-LLMs
Encoder- only Model Time #Parameters Base Model Pretraining Dataset Capability Data Type Open- Source
DNABERT (Ji et al., 2021) 2021.02 110M BERT GRCh38 Func. pred. DNA
Enformer (Avsec et al., 2021) 2021.10 240M BERT GRCh38 Func. pred. DNA
MoDNA (An et al., 2022) 2022.08 - BERT GRCh38 Func. pred. DNA $\times$
GPN (Benegas et al., 2023b) 2022.08 - Conv NCBI-Genome Variant pred. DNA
iEnhancer-BERT (Luo et al., 2022a) 2022.08 110M BERT - Enhancer. pred. DNA $\times$
iDNA-ABF (Jin et al., 2022) 2022.10 110M BERT - Func. pred. DNA
iEnhancer-ELM (Li et al., 2023e) 2022.12 110M BERT - Enhancer. pred. DNA
SA DNA-LM (Gankin et al., 2023) 2023.01 - BERT Ensembl Expression. pred. DNA
Nucleotide- Transformer (Dalla-Torre et al., 2023) 2023.01 50M-2.5B BERT 1000 Genomes Project, etc. Func. pred. DNA
SpliceBERT (Chen et al., 2023c) 2023.02 19.4M BERT UCSC Splicing pred. RNA
miProBERT (Wang et al., 2023a) 2023.03 110M BERT EPDnew Promoter pred. RNA $\times$
RNA-MSM (Zhang et al., 2023b) 2023.03 - BERT RNAcmap Structure pred. RNA
TFBert (Luo et al., 2023a) 2023.05 - BERT 690 ChIP-seq Binding pred. DNA
MRM-BERT (Zhang et al., 2023a) 2023.06 110M BERT - Modif. pred. RNA
GENA-LM (Fishman et al., 2023) 2023.06 110M-360M BERT GRCh38 Func. pred. and profiling DNA
DNABERT-2 (Zhou et al., 2023b) 2023.06 117M BERT GRCh38 Func. pred. DNA
PLPMpro (Li et al., 2023b) 2023.07 - BERT EPDnew Promoter pred. DNA $\times$
EpiGePT (Gao et al., 2023a) 2023.07 - BERT ENCODE Interpret genome DNA
Uni-RNA (Wang et al., 2023b) 2023.07 - BERT NCBI-Genome Func. pred. DNA $\times$
GPN-MSA (Benegas et al., 2023a) 2023.10 86M BERT - Variant class. and pred. DNA $\times$
Decoder- only GenSLMs (Zvyagin et al., 2022) 2022.10 25M-25B Diffusion+ GPT BV-BRC Evolution analysis RNA
HyenaDNA (Nguyen et al., 2023) 2023.06 0.44M-6.6M LongConv GRCh38 Func. pred. and class. DNA
DNAGPT (Zhang et al., 2023e) 2023.08 100M-3B GPT GRCh38 Func. pred. and generation DNA
Encoder- Decoder ENBED (Malusare et al., 2023) 2023.11 580M-1.2B Transformer NCBI-Genome Func. pred. and analysis DNA $\times$
Encoder-only Models. In the realm of encoder-only architecture, significant models such as SpliceBERT (Zheng et al., 2019), DNABERT (Ji et al., 2021), DNABERT-2 (Zhou et al., 2023b), iEnhancer-BERT (Luo et al., 2022a), scBERT (Yang et al., 2022), BioSeq-BLM (Li et al., 2021), PLPMpro (Li et al., 2023b), GPN (Benegas et al., 2022), SA DNA-LM (Gankin et al., 2023), RNA-MSM (Zhang et al., 2023b), miProBERT (Wang et al., 2023a), TFBert (Luo et al., 2023a), MRM-BERT (Zhang et al., 2023a), GPN-MSA (Benegas et al., 2023a), and iDNA-ABF (Jin et al., 2022) are noteworthy attempts employing the Transformer encoder within genomics. These models utilize a mask training mechanism where partial gene sequences are masked, prompting the model to predict and complete them, progressively learning inherent patterns within gene sequences. They demonstrate versatility in fine-tuning for various gene sequence-related tasks like promoter prediction, transcription factor binding site prediction, and single nucleotide variability, achieving commendable performance. MoDNA (An et al., 2022) employs a Bert-like encoder-only architecture, introducing a unique training paradigm with a stacked Generator-Discriminator structure. During training, the Generator completes masked sections based on sequence and motif information, while the Discriminator assesses the completed sequences. This dual-layer architecture enables the simultaneous integration of sequence features and motif information, facilitating motif-oriented learning. GENA-LM (Fishman et al., 2023) introduces a suite of encoder-based foundational DNA language models designed for genomics research. These models, unique in their ability to handle long sequences of up to 36,000 base pairs, address the challenge of decoding genomic sequences, which requires understanding extensive contextual information across numerous nucleotides. The integration of a newly-developed recurrent memory mechanism allows these models to process even larger DNA segments. Nucleotide-Transformer (Dalla-Torre et al., 2023) is a foundation encoder-only model pre-trained on DNA sequences. It integrates data from 3,202 diverse human genomes and 850 genomes from various species. The study focuses on bridging the gap in predicting molecular phenotypes from DNA sequences, a task often limited by the scarcity of annotated data and difficulties in transferring learnings between different prediction tasks. These transformer models, with parameters ranging from 50M to 2.5B, provide transferable, context-specific representations of nucleotide sequences, enabling more accurate molecular phenotype predictions, even in scenarios with limited data. Remarkably, these models learn to focus on key genomic elements, such as enhancers that regulate gene expression, without supervision, and can help prioritize functional genetic variants. EpiGePT (Gao et al., 2023a) is a transformer-based language pretrained model designed for epigenomics. It predicts genome-wide epigenomic signals by considering transcriptional regulation mechanisms. EpiGePT uniquely incorporates context-specific activities of transcription factors, offering deeper insights into gene regulation compared to models trained only on DNA sequences. It shows exceptional performance in diverse epigenomic signal prediction tasks and can learn cell-type-specific long-range interactions through self-attention mechanisms. Uni-RNA (Wang et al., 2023b) is an encoder-only model that excels in predicting RNA structures and functions, outperforming previous methods in various tasks. It leverages a billion RNA sequences to extract hidden evolutionary and structural information, revolutionizing RNA research and facilitating drug development. As gene sequences often significantly exceed the length of natural language text, many models aim to overcome the quadratic time complexity of attention mechanisms. Models like Enformer (Avsec et al., 2021) and LOGO (Yang et al., 2021) employ a blend of convolutional downsampling to handle longer DNA sequences efficiently. Tools like BioSeq-BLM (Li et al., 2021) integrate traditional analysis methods with language models. PLPMpro (Li et al., 2023b) marks the advent of pre-training and fine-tuning in this domain, comparing language models’ effects with traditional methods.
Decoder-only Models. Decoder-only models have gained significant attention due to their potent generative capacity. GenSLMs (Zvyagin et al., 2022) leverage genome-scale language models comprising multiple layers of attention-based decoders to elucidate the evolutionary dynamics of SARS-CoV-2. GenSLMs thus adeptly capture the evolutionary landscape of the SARS-CoV-2 genome. Pre-trained on prokaryotic gene sequences and further fine-tuned for SARS-CoV-2 genomes, GenSLM demonstrates precise and rapid identification of concerning variants. Besides commonly used architectures like transformers, niche architectures also find applications in genomic research. For instance, Models like DNAGPT (Zhang et al., 2023e) have introduced decoder-only architectures, akin to GPT, into the genomics domain by integrating specific symbols to guide different tasks, enabling zero-shot capabilities in species identification and regulatory factor prediction. HyenaDNA (Nguyen et al., 2023) incorporates a fast Fourier transform-accelerated recursive long-range convolution, a quadratic time complexity solution, enabling the handling of ultra-long DNA sequences. Employing convolution on the entire DNA sequence preserves single-nucleotide resolution, facilitating research into single nucleotide variability.
Encoder-Decoder Models. Encoder-decoder models in genomics, such as the Ensemble Nucleotide Byte-level Encoder-Decoder (ENBED) (Malusare et al., 2023), represent a significant advancement in bioinformatics. These models combine the strengths of both encoder and decoder components to effectively analyze and interpret complex genomic sequences. The encoder part of the model is responsible for compressing and encoding the input genomic data into a meaningful representation, capturing the essential features and patterns within the DNA sequences. This encoded representation is then passed to the decoder part, which is tasked with generating or reconstructing sequences, making predictions, or performing other relevant bioinformatics tasks.
6.2. Datasets
Table 8. Summary of datasets for Gene-LLMs
Dataset Time Scale Keywords
Pre-training Mammalian-Gene (Strausberg et al., 1999) 1999.10 - Mammalian Genomic Data
GRCh38 (Guo et al., 2017) 2013.12 3.2G nucleotides Genome Reference, Human, Annotations
690 ChIP-seq (Zeng et al., 2016) 2016.06 - DNA-binding Protein Data
DeepSEA (Zhou and Troyanskaya, 2015) 2017.04 - Chromatin Effects
1000 Genomes Project (Consortium et al., 2015) 2017.10 20.5T nucleotides Human Genetic Variation
EPDnew (Dreos et al., 2013) 2019.11 187K promoters Promoter Prediction
Panglao (Franzén et al., 2019) 2020.03 4M cells scRNA-seq Data
ExPecto (Rusk, 2018) 2020.12 - Gene Expression
UCSC-Genome (Karolchik et al., 2003) 2022.11 - Various Bio Data
BV-BRC (Olson et al., 2023) 2023.01 - Bacterial and Viral Pathogens
Ensembl (Lamesch et al., 2012) 2023.02 - Various Bio Data
RNAcmap (Rusk, 2018) 2023.07 - RNA Evolutionary analysis
ENCODE (coordination, data analysis coordination) 2023.09 - Functional Genomics
NCBI-Genome 2023.10 - Various Bio Data
TAIR (Lamesch et al., 2012) 2023.12 33602 genes Arabidopsis Information
VGDB (Hiscock and Upton, 2000) 2023.12 2847 genes Viral Genome Sequence
CAGI5 (Katsonis and Lichtarge, 2019) 2019.07 - Assessing Genetic and Genomic Predictions
Benchmark Protein-RNA (Horlacher et al., 2023) 2023.08 - Evaluating RBP-RNA Interaction Prediction Methods
NT-Bench (Dalla-Torre et al., 2023) 2023.09 3.2G+20.5T+174G nucleotides Evaluating Genomics Foundational Models
Currently, the majority of Gene-LLMs are trained on human DNA data. To increase the amount of training data and enhance the model’s generalization ability, genomes of organisms closely related to humans, such as chimpanzees or pigs, or even the genomes of all mammals are utilized for training. Here we list prevalent training datasets and benchmarks for Gene-LLMs, as summarized in Table 8.
Pre-training Datasets:
  • Mammalian-Gene (Eppig et al., 2017). In Gene-LLM training, integrating mammalian genomic data alongside human data enhances the model’s generalization. These databases provide a plethora of genomic sequences from various mammals, aiding in building a robust Gene-LLM capable of better understanding and interpreting genomic sequences across different species, thereby enriching genomic research.
  • GRCh38 (Human Genome) (Guo et al., 2017). Human genome is frequently used as the training corpus for Gene-LLMs. Genome Reference Consortium Human Build 38 (GRCh38) is a substantial assembly of the human genome released after the previous major assembly. The genomic sequences, annotations, and metadata contained within GRCh38 provide a rich source of information that can be utilized to train LLMs to recognize various genomic features, understand genomic syntax, and make predictions related to gene expression, mutation impacts, and other genomic phenomena.
  • 690 ChIP-seq (Zeng et al., 2016) The 690 ChIP-seq datasets are a comprehensive collection of DNA-binding protein data, encompassing 91 human cell types and 161 specific DNA-binding proteins under various conditions. The significance of these datasets lies in their diversity and volume, providing a rich foundation for understanding DNA-protein interactions. This variety and scale are crucial for training advanced models like GeneLLM, as they allow the model to learn from a wide range of scenarios, ensuring robustness and generalizability in predicting DNA-protein bindings. The inclusion of data from different cell types and conditions also aids in capturing the complexity of gene regulation, enhancing the model’s ability to make accurate predictions in various biological contexts.
  • DeepSEA (Zhou and Troyanskaya, 2015). DeepSEA compiles extensive chromatin profiles from ENCODE and Roadmap Epigenomics projects, comprising data on 690 transcription factor bindings, 125 DNase I sensitivity profiles, and 104 histone-mark profiles. It is a significant advance for researchers looking to prioritize functional variants, including eQTLs and disease-associated variants, using purely genomic sequence data, thus offering a robust resource for gene model training and functional genomic studies.
  • 1000 Genomes Project (Consortium et al., 2015) The 1000 Genomes Project provided a comprehensive reference for human genetic variation. It sequenced over 2,500 individuals from 26 populations, using whole-genome sequencing, deep exome sequencing, and microarray genotyping. The study identified 88 million variants, including single nucleotide polymorphisms, insertions/deletions, and structural variants. This work expanded our understanding of genetic diversity, particularly in underrepresented populations, and has been instrumental in studies of human evolution, disease biology, and genetic research methodologies. The data serves as a dataset for human genetic studies and has greatly enhanced the catalog of genetic variation, particularly in South Asian and African populations.
  • EPDnew (Dreos et al., 2013). This is a collection of databases encompassing experimentally validated promoters for selected model organisms, with evidence derived from TSS-mapping from high-throughput experiments like CAGE and Oligo-capping. This data can be particularly useful for tasks related to promoter prediction, gene expression analysis, and other genomic sequence interpretation tasks within the realm of genomics and bioinformatics.
  • Panglao (Franzén et al., 2019) PanglaoDB is a comprehensive web server database designed for the exploration of mouse and human single-cell RNA sequencing (scRNA-seq) data. This database stands out due to its user-friendly interface, allowing easy access and exploration of extensive scRNA-seq datasets. With over 1054 single-cell experiments, encompassing more than 4 million cells from a diverse range of tissues and organs, PanglaoDB presents an invaluable resource for biological research. The availability of pre-processed and quality-controlled scRNA-seq data allows for efficient training and validation of machine learning models aimed at gene expression analysis. The database’s broad coverage of cell types and conditions provides a diverse training set that can help improve the accuracy and generalizability of predictive models in genomics.
  • ExPecto (Rusk, 2018) The ExPecto dataset is designed for predicting tissue-specific gene expression from DNA sequences. This dataset was validated using known expression quantitative trait loci and massively parallel reporter assays, and it has been used to prioritize variants in genome-wide association studies. It facilitates the characterization of genetic variation potential within each gene and enables systematic investigation into tissue-specific effects of mutations on gene expression.
  • UCSC-Genome (Karolchik et al., 2003). The UCSC Genome Database is a comprehensive online tool for genomic research. It provides access to genomic sequences from various species, integrated with a plethora of aligned annotations. This browser facilitates rapid visualization and analysis of genomic data, supported by an extensive database and a suite of interactive tools. With over 108 species’ genomes, it allows for in-depth comparative genomics and includes unique features like Assembly Hubs and Custom Tracks, enabling researchers to personalize data analysis.
  • BV-BRC (Olson et al., 2023). The BV-BRC dataset is a comprehensive resource integrating genomic and other related data for bacterial and viral pathogens. It hosts a vast collection of genomic sequences, including over 600,000 bacterial genomes, 11,000 archaeal genomes, and 8.5 million viral genomes, which encompasses over 6 million SARS-CoV-2 genomes. This dataset is instrumental in comparing human pathogens with their non-pathogenic relatives, thereby aiding in understanding virulence and pathogenicity. The BV-BRC also includes metadata and non-genomic data like protein structures, enhancing research in host-pathogen interactions and responses to infections. The data is consistently annotated, and its extensive nature supports a wide range of comparative genomic analyses and epidemiological studies.
  • Ensembl (Martin et al., 2023). It offers centralized access to a wide range of data, including DNA sequences, gene models, annotations, and variation data for vertebrates and model organisms. This platform is instrumental for those studying genomics and serves as a repository for large-scale data suited for training complex models. Ensembl’s automated annotation and vast database make it an invaluable resource for machine learning and bioinformatics, providing the raw material for the development of predictive models and analysis tools.
  • RNAcmap (Rusk, 2018). The RNAcmap dataset focuses on RNA evolutionary coupling analysis. It employs a fully automatic pipeline, RNAcmap, which doesn’t rely on pre-curated Rfam families. This pipeline uses BLAST-N for initial homolog searches and INFERNAL for building a covariance model based on predicted secondary structures. The dataset supports the generation of base-pairing and distance-based contact maps, using multiple sequence alignment from secondary searches. It demonstrates the less dependency on specific evolutionary coupling tools and more on the accuracy of secondary structure predictors. This dataset is beneficial for RNA structure prediction, offering a significant step in computational RNA analysis by enabling evolutionary coupling analysis for any RNA sequence.
  • ENCODE (Feingold et al., 2004). The Encyclopedia of DNA Elements (ENCODE) Project aimed at developing a comprehensive map of functional elements within the human genome, which includes genes among other functional genomic elements. This rich dataset offers a vast amount of functional genomics data, which encompasses a variety of experiments and annotations related to stem cell development, immune cells, mouse development, and other crucial biological processes.
  • NCBI-Genome [^10] https://www.ncbi.nlm.nih.gov. The National Center for Biotechnology Information (NCBI) is a key platform offering comprehensive data for gene-related research and large model training. NCBI hosts a multitude of databases such as GenBank for DNA sequences, which is crucial for understanding genetic blueprints across various organisms. Their Gene database provides a vital repository for gene-specific information, linking genomic maps, sequences, expression, and function data, making it an invaluable resource for bioinformatics.
  • TAIR (Lamesch et al., 2012). Arabidopsis Information Resource (TAIR) is a comprehensive database for the model plant Arabidopsis thaliana, offering a wealth of data for large-scale model training. It provides annotated genomic sequences, gene function and expression data, metabolic pathways, and gene expression patterns. The database integrates RNA-seq and proteomics data, offering a rich set of annotations for molecular function, cellular components, and biological process aspects of the GO terms.
  • VGDB (Hiscock and Upton, 2000). Viral Genome DataBase is a substantial resource organizing the ‘sequence space’ of viral genomes in a structured manner, aiming to serve as an annotated and curated database of complete viral genome sequences along with value-added derived data and data mining tools. Moreover, VGDB offers detailed gene information and predicted protein sequences from completely sequenced genomes of large viruses, encompassing DNA sequence data, protein sequence data, and other related annotations.
Benchmarks:
  • CAGI5 Challenge Benchmark (CAGI5) (Cheng et al., 2019). The CAGI5 Challenge is a comprehensive benchmark designed to rigorously assess computational methods in predicting a wide array of genetic and genomic outcomes. At the forefront is the prediction of effects of 17,500 single nucleotide variants and small indels in 11 disease-associated enhancers and 10 promoters using a saturation mutagenesis massively parallel reporter assay. Additionally, the challenge delves into the functional effects of nonsynonymous variants in proteins. Another significant aspect of the challenge is the prediction of the impact of variants on exon splicing. The challenge also extends to clinical genomes, where participants are tasked with matching patients’ genome sequences to their clinical descriptions and predicting the causal pathogenic variants. This includes data from children with suspected genetic disorders and gene panel sequences associated with intellectual disability or Autism spectrum disorders.
  • Protein-RNA Interaction Prediction Benchmark (Protein-RNA) (Horlacher et al., 2023). While numerous machine-learning-based methods have been developed to predict RNA-binding protein (RBP) interactions, direct comparisons of their performance have been challenging due to the use of varied training and evaluation datasets across different RBP sets and CLIP-seq protocols. It compiled a set of 37 methods designed for in vivo RBP-RNA interaction prediction. This benchmark systematically evaluates the performance of 11 RBP binding-site prediction methods using binding sites from three large CLIP-seq repositories, which include 313 unique CLIP-seq experiments across 203 RBPs.
  • Nucleotide Transformer Benchmark (NT-Bench) (Dalla-Torre et al., 2023). This is a comprehensive evaluation framework designed to assess the performance of genomics foundational models. This benchmark pits the Nucleotide Transformer models against other prominent genomics models, such as DNABERT, HyenaDNA (with both 1kb and 32kb context lengths), and Enformer. To ensure a level playing field, all models in this benchmark are subjected to fine-tuning and evaluation using a uniform protocol that spans 18 downstream tasks with the corresponding datasets (e.g., GRCh38, 1000 Genomes project, and the multispecies dataset from NCBI).
6.3. Evaluation
Gene-LLMs are primarily evaluated based on their capabilities in four critical areas: function prediction, structure prediction, sequence generation, and sequence variation and evolution analysis.
Function Prediction. The prediction of gene function has long been a pivotal focus in genomics research. Traditionally, models were trained directly on specific sequences relevant to these tasks to acquire the requisite capabilities. The advent of LLMs has revolutionized this approach by facilitating pre-training on vast amounts of genomic data, followed by targeted fine-tuning on specific tasks, empowering these models to effectively discern and predict gene expression patterns and their regulatory elements with enhanced accuracy and broader context. Here we present three subtasks that are frequently used for function prediction.
  • Promoter Prediction. One crucial task in genomics research involves predicting the locations of promoters, specific DNA segments recognized and initiated for transcription by RNA polymerase. Promoters, typically found upstream of structural genes, contain essential sequences vital for RNA polymerase binding and transcription initiation. DNABERT (Ji et al., 2021), for instance, effectively utilized human TATA and non-TATA promoter regions to fine-tune downstream models, surpassing existing state-of-the-art models.
  • Enhancer Prediction. Enhancers, short DNA regions, play a pivotal role in enhancing the likelihood of gene transcription by providing binding sites for various molecules, facilitating the assembly and stabilization of transcription machinery. Accurate prediction of enhancers holds significant importance in comprehending gene regulation, a critical aspect for understanding health and disease. Identifying enhancers can unveil genetic foundations of diseases and offer potential targets for therapeutic interventions. Notably, encoder-only models like iEnhancer-BERT (Luo et al., 2022a), iEnhancer-ELM (Li et al., 2023e) basically follow the same method, i.e., pretraining on human genome sequences and fintuning on enhancer data.
  • Binding Site Prediction. Binding site prediction aims to pinpoint regions on DNA, RNA, or proteins where specific molecules, such as transcription factors or peptides, can bind. EpiGePT (Gao et al., 2023a) employs a bin-scanning strategy to detect potential binding sites for a vast set of human transcription factors, selecting the maximum score for each factor. By integrating motif and expression features, EpiGePT offers a comprehensive representation for each DNA bin, significantly enhancing the accuracy of binding site predictions.
Structure Prediction. Structure Prediction involves computational tools for identifying and modeling biologically significant nucleic acid structures. These tools also aid in designing novel molecular architectures for nanotechnology and synthetic biology. The field has seen progress in predicting RNA three-dimensional structure directly from sequence and in designing sequences for predefined DNA and RNA nanostructures, indicating that nucleic acid structure is both predictable and controllable. Here we present two subtasks that are frequently used for structure prediction.
  • Chromatin Profile Prediction. It involves the accurate prediction of the intricate structure of chromatin—a complex formed by DNA and proteins within cells. Understanding these predictions unveils crucial insights into gene regulation mechanisms and the role of chromatin alterations in diseases, notably cancer. HyenaDNA (Nguyen et al., 2023) concentrates on predicting chromatin profiles and epigenetic markers from DNA sequences, quantifying functional effects caused by non-coding variants.
  • DNA/RNA-Protein Interaction Prediction. It aims to understand the intricate interplay between nucleic acids and proteins. This interaction is pivotal for numerous biological processes, including transcriptional regulation, RNA splicing, and translation. TFBert (Luo et al., 2023a) leverages the power of pre-trained BERT architectures for predicting DNA-protein interactions.
Sequence Generation. Generation ability stands as a pivotal task in bioinformatics, focusing on models’ capacity to create artificial sequences that closely mimic real biological sequences. This ability becomes particularly crucial when considering the generation of artificial human genomes (AGs), serving as tools to safeguard genetic privacy and reduce costs linked with genetic sample collection. DNAGPT (Zhang et al., 2023e) emerges as a model purpose-built for this task. This model generated 5,000 AGs, each spanning a region of 10,000 Single Nucleotide Polymorphisms (SNPs). When compared with other models, DNAGPT showcased superior performance across various metrics.
Sequence Variation and Evolution Analysis. It plays a pivotal role in deciphering DNA sequence alterations and their evolutionary trajectories. Understanding these variations is essential in unraveling the genetic foundations of traits, diseases, and evolutionary patterns. GenSLMs (Zvyagin et al., 2022) were adapted from large language models to analyze genomic data, specifically focusing on the evolutionary landscape of SARS-CoV-2 genomes. In a different vein, the GPN-MSA (Benegas et al., 2023a) model introduces a novel framework for DNA language models by harnessing whole-genome sequence alignments across multiple species.
Evaluation Metric. The evaluation metrics for function prediction mainly include accuracy, F1-score, sensitivity, specificity, and precision. For structure prediction, a common evaluation approach is to assess the accuracy of predicted DNA structures against known structures. Median AUROC (Area Under the Receiver Operating Characteristic Curve) is often used for chromatin profile prediction in structure prediction. In sequence generation tasks, the evaluation of generated labels plays a pivotal role. Metrics like Matthews correlation coefficient (MCC), perpendicular distance, true positive rate, and false positive rate values are widely used for label classification. As for the sequence variation and evolution analysis, AUROC can be used for classifying pathogenic variants. Additionally, the accuracy in identifying variants of concern is also a critical metric.
6.4. Summary
This section provides a comprehensive overview of Genomic Large Language Models (Gene-LLMs), detailing their application in computational biology for analyzing DNA and RNA sequences. It categorizes Gene-LLMs into encoder-only, decoder-only, and encoder-decoder models, highlighting their architecture, datasets, and evaluation methods. The discussion emphasizes the significance of these models in understanding gene function, chromatin profile prediction, splice site prediction, binding site prediction, and structure prediction, showcasing their capabilities in generating sequences and analyzing sequence variation and evolution. However, it is evident that the availability of generative Gene-LLMs (e.g., decode-based models) is severely limited, necessitating the development of robust generative models to facilitate diverse genomics applications.
7. Multi-modal Scientific Large Language Models
Multimodal Large Language Models have emerged as a prominent research area, utilizing robust LLMs as the core to handle multimodal data. These models possess a distinctive capability to incorporate diverse data types like text, images, audio, and other forms of information, enabling comprehensive exploration and problem-solving across various domains. The multimodal models hold promising prospects, particularly in the realms of biological and chemical sciences encompassing protein, molecular, and genomic studies. In this section, we explore recent advancements in multimodal models within these scientific fields (i.e., MM-Sci-LLMs), highlighting their capabilities and the exploited datasets. Note that, this survey focuses on cross-lingual multimodal models, involving at least two languages from different domains, such as text and molecule. Hence, we exclude the monolingual multimodal approaches from MM-Sci-LLMs, like the joint modeling of protein sequences and structures (Su et al., 2023; Wang et al., 2023d; Zhang et al., 2023c). The diverse forms of different languages and modalities are illustrated in Figure 4, and the overview of this section is shown in Figure 11.
{forest}
for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw, rounded corners,align=left, minimum width=2.5em, inner xsep=4pt, inner ysep=1pt, , where level=1text width=5em,fill=blue!10, where level=2text width=5em,font=,fill=pink!30, where level=3font=,yshift=0.26pt,fill=yellow!20, [MM-Sci-LLMs, fill=green!10 [Models
(Sec. 7.1),text width=4em [Molecule-Text, text width=5.5em [Encoder-only: Text2Mol (Edwards et al., 2021)   KV-PLM (Zeng et al., 2022)  MoMu (Zeng et al., 2022)  MolFM (Luo et al., 2023b) GPT-MolBERTa (Balaji et al., 2023) ] [Decoder-only: DrugChat (Liang et al., 2023)  MolReGPT (Li et al., 2023d)  MolXPT (Liu et al., 2023e) ] [Encoder-Decoder: MolT5 (Edwards et al., 2022)  Text+Chem T5 (Edwards et al., 2022)  ChatMol (Zeng et al., 2023b)  GIT-Mol (Liu et al., 2023b)
MoleculeSTM (Liu et al., 2023a) ] ] [Protein-Text, text width=5.5em [Encoder-only: ProTranslator (Xu and Wang, 2022)  ProtST-ProtBert (Xu et al., 2023b) ] [Decoder-only: InstructProtein (Wang et al., 2023c) ] [Encoder-Decoder: ProteinDT (Liu et al., 2023f)  Prot2Text (Abdine et al., 2023)  ProtST-ESM-1B (Xu et al., 2023b)  ProtST-ESM-2 (Xu et al., 2023b) ] ] [Protein-Molecule, text width=5.5em [Encoder-only: DrugCLIP (Gao et al., 2023b)] [Decoder-only: DrugGPT (Li et al., 2023a) ] [Encoder-Decoder: ChemBERTaLM (Uludoğan et al., 2022)  DeepTarget (Chen et al., 2023b) ] ] [Comprehensive, text width=5.5em [Encoder-only:BioTranslator (Xu et al., 2023a)] [Decoder-only: Galactica (Taylor et al., 2022b)  ChatDrug (Liu et al., 2023c)  DARWIN-MDP (Xie et al., 2023)
BioMedGPT-10B (Luo et al., 2023c) Mol-Instructions (Fang et al., 2023a) ] [Encoder-Decoder: BioT5 (Pei et al., 2023) ] ] ] [Datasets
(Sec. 7.2),text width=4em [Training, text width=3.5em [SwissProtCLAP (Liu et al., 2023f)   ProtDescribe (Xu et al., 2023b)  InstructProtein (Wang et al., 2023c)  Scientific Knowledge Dataset (Xie et al., 2023)
BioLip (Yang et al., 2012)  ChEBL-dia (Zeng et al., 2023b)  PubChemQA (Luo et al., 2023c)  Prot2Text (Abdine et al., 2023)  PCdes (Zeng et al., 2022)  PubChemSTM (Liu et al., 2023a)
MoMu (Zeng et al., 2022)  ChEBI-20 (Edwards et al., 2021)  Mol-Instructions (Fang et al., 2023a) ] ] [Benchmark, text width=3.5em [ MoleculeNet  (Wu et al., 2018)   BindingDB (Gilson et al., 2016)   DUD-E (Mysinger et al., 2012)  ] ] ] [Evaluation
(Sec. 7.3),text width=4em [Cross-modal Prediction, text width=7.5em [Text-based Function prediction across different modalities ] ] [Cross-modal Retrieval, text width=7.5em [Mutual retrieval across different modalities ] ] [Cross-modal Generation, text width=7.5em [Mutual generation across different modalities ] ] [Evaluation Metric, text width=7.5em [ Refer to Sec. 3.3; 4.3; 5.3; 6.3 ] ] ] [Summary
(Sec. 7.4),text width=4em ] ]
Figure 11. Chapter overview of MM-Sci-LLMs.
Table 9. Summary of MM-Sci-LLMs
Model Time #Parameters Base Model Pretraining Dataset Open- source
Mol.&Text Text2Mol (Edwards et al., 2021) 2021.11 117M GCN, SciBERT PubChem, ChEBI
KV-PLM (Zeng et al., 2022) 2022.02 110M SciBERT MoleculeNet, USPTO 1k TPL (Schneider et al., 2016), Chemprot (Kringelum et al., 2016), BC5CDR (Li et al., 2016)
MoMu (Su et al., 2022) 2022.09 110M SciBERT, GIN S2ORC (Lo et al., 2020), PubChem, OGB (Hu et al., 2020)
MolT5 (Edwards et al., 2022) 2022.11 60M T5 ChEBI-20
Text+Chem T5 (Christofidellis et al., 2023) 2023.05 40M/220M T5 ChEBI-20
DrugChat (Liang et al., 2023) 2023.05 13B GNN, Vicuna PubChem, ChEMBL
MolReGPT (Li et al., 2023d) 2023.06 - ChatGPT ChEBI-20
ChatMol (Zeng et al., 2023b) 2023.06 - T5 ChEBL-dia (Zeng et al., 2023b)
MolXPT (Liu et al., 2023e) 2023.05 350M GPT-2 PubChem, PubMed $\times$
MolFM (Luo et al., 2023b) 2023.07 138M GIN, Trans. enc., TransE (Bordes et al., 2013) PubChem, S2ORC
GIT-Mol (Liu et al., 2023b) 2023.08 700M SciBERT, GIN, SwinTransformer (Liu et al., 2021a) PubChem, ChEBI-20 $\times$
GPT-MolBERTa (Balaji et al., 2023) 2023.10 - RoBERTa MoleculeNet
MoleculeSTM (Liu et al., 2023a) 2023.12 252M GraphMVP (Liu et al., 2021b), SciBERT Chemformer (Irwin et al., 2022) PubChemSTM
Prot.&Text ProTranslator (Xu and Wang, 2022) 2022.04 - PubMedBert -
ProteinDT (Liu et al., 2023f) 2023.02 - SciBERT, ProtBERT SwissProtCLAP $\times$
ProtST-ProtBert (Xu et al., 2023b) 2023.07 420M ProtBert, PubMedBERT ProtDescribe
ProtST-ESM-1b (Xu et al., 2023b) 2023.07 650M ESM-1b, PubMedBERT ProtDescribe
ProtST-ESM-2 (Xu et al., 2023b) 2023.07 650M ESM-2, PubMedBERT ProtDescribe
Prot2Text (Abdine et al., 2023) 2023.07 283M ESM2, RGCN (Schlichtkrull et al., 2018), Trans. dec. Prot2Text $\times$
InstructProtein (Wang et al., 2023c) 2023.10 1.3B OPT (Zhang et al., 2022a) InstructProtein $\times$
Mol.&Prot. ChemBERTaLM (Uludoğan et al., 2022) 2022.09 - RoBERTa, ChemBERTa BindingDB, MOSES (Polykovskiy et al., 2020)
DeepTarget (Chen et al., 2023b) 2023.03 - Transformer ChEMBL
DrugGPT (Li et al., 2023a) 2023.06 1.5B GPT-2 ZINC20
DrugCLIP (Gao et al., 2023b) 2023.10 - SE(3) 3D Transformer PDBBind (Wang et al., 2005), BioLip, ChEMBL $\times$
Comprehensive Galactica (Taylor et al., 2022b) 2022.11 120B Trans. dec. -
BioTranslator (Xu et al., 2023a) 2023.02 - PubMedBERT -
ChatDrug (Liu et al., 2023c) 2023.05 - ChatGPT -
BioMedGPT-10B (Luo et al., 2023c) 2023.08 10B LLaMA2, GIN, ESM-2 PubChemQA, UniProtQA
DARWIN-MDP (Xie et al., 2023) 2023.08 7B LLaMA SciQ, FAIR (Wilkinson et al., 2016)
BioT5 (Pei et al., 2023) 2023.10 252M T5 ZINC20
Mol-Instructions (Fang et al., 2023a) 2023.11 - LLaMA Mol-Instructions
7.1. Models
To provide a comprehensive overview of MM-Sci-LLMs, we categorize them into four groups: molecule-text models, protein-text models, protein-molecule models, and comprehensive models, based on the specific modality they focus on. Table 9 shows the summary of MM-Sci-LLMs.
7.1.1. Molecule-Text Models
Molecule-text models are designed to understand both the chemical structure and text descriptions of molecules by creating associations between text-molecular structure data pairs. This approach enables the model to effectively discern distinctions between various molecules, enhancing its ability for generalization.
Text2Mol (Edwards et al., 2021), KV-PLM (Zeng et al., 2022), MoMu (Su et al., 2022), MolFM (Luo et al., 2023b), GPT-MolBERTa (Balaji et al., 2023) and MoleculeSTM  (Liu et al., 2023a) represent a suite of encoder-based molecule-text multimodal models. Text2Mol (Edwards et al., 2021) is a multimodal embedding approach to establish an aligned semantic space between text and molecules, facilitating cross-modal retrieval. This approach introduces the CHEBI-20 dataset and employs cross-modal attention-based association rules to enhance the performance in downstream tasks. KV-PLM (Zeng et al., 2022) represents a cross-modal learning approach that integrates molecular structural information with SMILES into textual knowledge. Similar to BERT’s pre-training in an unsupervised fashion, this method facilitates the acquisition of meta-knowledge regarding structural fragments within the textual context. MoMu (Su et al., 2022) is a molecular multimodal model pretrained on both molecular graphs with Graph Isomorphism Network (GIN) (Xu et al., 2018) and their semantically related textual data with SciBERT (Beltagy et al., 2019) via contrastive learning, establishing a crucial link between molecular graphs and natural language. MolFM (Luo et al., 2023b) serves as a multimodal molecular foundation model, facilitating joint representation learning from molecular structures, descriptive texts, and knowledge graphs. It proposes a cross-modal attention mechanism that enhances the comprehension of molecular structures by incorporating atoms, neighboring molecule entities, and semantically related texts. GPT-MolBERTa (Balaji et al., 2023) utilizes ChatGPT to generate rich textual descriptions of molecules, which is then used to pre-train the RoBERTa model in a self-supervised manner. MoleculeSTM (Liu et al., 2023a) employs a molecule encoder and a text encoder with a contrastive learning strategy to jointly learn the chemical structures and their textual descriptions.
DrugChat (Liang et al., 2023), MolXPT (Liu et al., 2023e) and MolReGPT (Li et al., 2023d) are decoder-based molecule-text multimodal models. DrugChat (Liang et al., 2023) combines a graph neural network (GNN) and Vicuna (Chiang et al., 2023) (a classical LLM) to enable ChatGPT-like capabilities on drug molecule graphs. MolXPT (Liu et al., 2023e) is a generative pre-trained model that jointly models text and molecules. It aims to enhance the interaction between molecules and text, facilitating tasks such as molecular property prediction and text-guided molecule generation. MolReGPT (Li et al., 2023d) framework utilizes in-context few-shot molecule learning to enable language models to perform molecule-caption translation tasks without fine-tuning. MolT5 (Edwards et al., 2022), Text+Chem T5 (Christofidellis et al., 2023), ChatMol (Zeng et al., 2023b) and GIT-Mol (Liu et al., 2023b) use encoder-decoder architectures. MolT5 (Edwards et al., 2022) is a T5-based self-supervised learning framework that allows for the translation between molecules and natural language. Text+Chem T5 (Christofidellis et al., 2023) leverages multi-task learning to unify molecular and textual representations, which can solve language tasks, chemical tasks, and cross-domain tasks, without the need for task-specific fine-tuning or retraining. ChatMol (Zeng et al., 2023b) is initialized with T5 and conducts multi-task pre-training, facilitating conversational molecular design through natural language interaction. Based on Q-Former design from BLIP2 (Li et al., 2023c), GIT-Mol (Liu et al., 2023b) proposes a novel modality mixer designed with a cross-attention mechanism, enabling seamless fusion of multiple molecular modalities, including SMILES, texts, images, and structure graphs.
7.1.2. Protein-Text Models
The utilization of textual data from literature and databases has become increasingly important in protein research, leading to the emergence of protein-text multimodal models. ProTranslator (Xu and Wang, 2022) innovatively approaches protein function prediction by redefining it as a machine translation problem. By utilizing textual descriptions, ProTranslator can annotate proteins to functions without the need for pre-known associated proteins. ProteinDT (Liu et al., 2023f) adopts contrastive learning to align the representation space of text and protein sequences. The ProteinDT model addresses the problem of aligning text and protein sequences to enable text-guided protein design and editing tasks. ProtST series (Xu et al., 2023b), including ProtST-ESM-1B, ProtST-ESM-2, and ProtST-ProtBert, are models enhancing protein sequence pre-training by integrating biomedical texts. Leveraging tasks like unimodal mask prediction and multimodal representation alignment, these models demonstrate superior performance in representation learning benchmarks. Prot2Text (Abdine et al., 2023) predicts a protein’s function in free text format by integrating GNNs and LLMs in an encoder-decoder framework. InstructProtein (Wang et al., 2023c) stands out as a decode-only LLM bridging the gap between human and protein languages through bidirectional generative capabilities.
7.1.3. Protein-Molecule Models
Protein-molecule models focus on converting protein information into molecular information and exploring potential associations between the two. Protein-molecular models typically employ an encoder-decoder architecture, where the encoder converts a protein sequence into fixed-length vectors, and the decoder generates molecular information from those vectors. ChemBERTaLM (Uludoğan et al., 2022) uses a pre-trained biochemical language model to initialize a target molecule generation model, aiming to address the challenge of limited data on new protein labeling in drug discovery. DeepTarget (Chen et al., 2023b) is an innovative deep generative model that designs new drug molecules directly from protein amino acid sequences, integrating adversarial networks and contrastive learning to minimize reliance on existing structural information. DrugGPT (Li et al., 2023a) is built upon the GPT framework for ligand design, enabling precise capture of both structural information and chemical principles governing drug molecules. DrugCLIP (Gao et al., 2023b) is a novel approach to virtual screening for potential drugs that uses contrastive learning to align representations of binding protein pockets and molecules. These methods aim to improve the efficiency and accuracy of drug discovery through AI-assisted methods.
7.1.4. Comprehensive Models
In this survey, the comprehensive multimodal models include text, proteins, molecules, genomes and other scientific languages. Their potential impact spans diverse domains, particularly offering significant support in accelerating fundamental science research. Galactica (Taylor et al., 2022b), BioTranslator (Xu et al., 2023a), ChatDrug (Liu et al., 2023c), BioMedGPT0-10B (Luo et al., 2023c), DARWIN-MDP  (Xie et al., 2023), BioT5 (Pei et al., 2023) and Mol-Instructions (Fang et al., 2023a) are representative comprehensive multimodal models. Galactica (Taylor et al., 2022b) is a scientific language model based on a decoder-only architecture, trained on a massive corpus of 106 billion tokens from various scientific texts and knowledge bases. Its strength lies in its ability to efficiently process extensive scientific datasets, enhancing its potential utility in specific scientific tasks. BioTranslator (Xu et al., 2023a) leverages PubMedBERT (Gu et al., 2021) for modeling textual descriptions and fine-tunes it using ontologies across various domains, showcasing its versatility in integrating diverse modalities for sophisticated biological data analysis. ChatDrug (Liu et al., 2023c) utilizes conversational language models to perform drug editing tasks. It addresses the problem of modifying small molecules, peptides, and proteins to achieve specific properties, such as solubility or secondary structure, by leveraging the knowledge extraction and summarization capabilities of the language model. BioMedGPT-10B (Luo et al., 2023c) is an open multimodal generative pre-trained transformer, to bridge the modality gap in biomedicine by aligning different biological modalities with natural language. It addresses the challenge of effectively communicating with diverse biological modalities through free text, including molecular structures, protein sequences, pathways, and cell transcriptomics. DARWIN-MDP (Xie et al., 2023) is a variant of the DARWIN series that involves 16 FAIR datasets to generate instructions for material and device prediction tasks, such as classification, regression, and design. BioT5 (Pei et al., 2023) is a pre-training framework that leverages structured and unstructured data sources to capture the underlying relations and properties of bio-entities. It aims to address the challenges in cross-modal generation and prediction tasks in the field of computational biology. Mol-Instructions (Fang et al., 2023a) aims to address the challenge of enhancing LLMs’ performance in tasks related to molecules, proteins, and biomolecular text by providing task-specific instructions.
7.2. Datasets
Table 10. Summary of datasets for MM-Sci-LLMs
Dataset Time Scale Keywords
Mol.&Text ChEBI-20 (Edwards et al., 2021) 2021.11 33K Cross-Lingual Retrieval
PCdes (Zeng et al., 2022) 2022.02 15K SMILES strings, Cross-information Retrieval
MoMu (Su et al., 2022) 2022.09 15K Bidirectional Molecular Graph-to-text Retrieval
PubChemSTM (Liu et al., 2023a) 2022.12 281K+250K Structure-Text Pairing
ChEBL-dia (Zeng et al., 2023b) 2023.06 10K Molecule Generation Evaluation
PubChemQA (Luo et al., 2023c) 2023.08 325K+365K Molecule Description
Prot.&Text SwissProtCLAP (Liu et al., 2023f) 2023.02 441K Protein Sequence Analysis
ProtDescribe (Xu et al., 2023b) 2023.07 553K Protein Sequence Representation
Prot2Text (Abdine et al., 2023) 2023.07 256K Protein Function Prediction
UniProtQA (Luo et al., 2023c) 2023.08 569K+1.8M Protein Function and Property
InstructProtein (Wang et al., 2023c) 2023.10 2.8M Structured KG, Bidirectional Protein-Text Generation
Mol.&Prot. DUD-E  (Mysinger et al., 2012) 2012.06 22k Ligand Diversity, molecular docking
BioLiP (Yang et al., 2012) 2012.10 204k ligand-protein interactions, molecular docking
BindingDB (Gilson et al., 2016) 2016.01 1.1M protein-ligand interactions, drug design
Comprehensive Galactica (Taylor et al., 2022b) 2022.11 62.2M Large Scientific Corpus
Scientific Knowledge Dataset (Xie et al., 2023) 2023.08 13k+6M biomolecular studies
Mol-Instructions (Fang et al., 2023a) 2023.10 2M Comprehensive Instruction, Diverse Text Descriptions
Similar to the taxonomy of multimodal models, we divide multimodal datasets into four categories. The following present popular datasets for each category, which is also summarized in Table 10.
7.2.1. Molecule-Text Datasets
The molecule-text datasets are mainly sourced from chemical literature, patents, drug instructions, etc., which contain descriptions of molecules. After pre-processing (e.g., data clearing), it can be used to train the model to achieve communication between molecular language and natural language. .
  • ChEBI-20 (Edwards et al., 2021). CHEBI-20 was created by collecting ChEBI (Hastings et al., 2016) annotations of compounds from PubChem (Kim et al., 2016, 2019), leading to a collection of 102,980 compound-description pairs. From this data, a subset of 33,010 pairs was curated to form the CHEBI-20 dataset. It can be applied to assess the model’s efficacy in molecular captioning tasks.
  • PCdes (Zeng et al., 2022). PCdes comprises SMILES notations and corresponding property descriptions for 15,000 molecules from PubChem (Wang et al., 2009). This dataset facilitates cross-information retrieval tasks. For instance, KV-PLM (Zeng et al., 2022), when fine-tuned on PCdes, aims to optimally select matching SMILES strings or property description sentences for each other.
  • MoMu (Su et al., 2022). MoMu was developed by aggregating SMILES strings from the first 50,000 molecular compounds in the PubChem (Wang et al., 2009) database and transforming them into molecular maps. This process was complemented by incorporating relevant scientific texts from the fields of medicine, biology, chemistry, and computer science, obtained from the S2ORC (Lo et al., 2020) database. The final dataset comprised 15,613 pairs of graphic documents. It can be used for many tasks and applications related to molecular structure and properties, including molecular image description generation, zero-sample molecular generation, molecular property prediction, and molecular similarity calculation.
  • PubChemSTM (Liu et al., 2023a). PubChemSTM dataset is constructed using data from the PubChem (Kim et al., 2021a) database, which consists of 250,000 molecules and 281,000 structure-text pairs. PubChemSTM is used to train the MoleculeSTM model (Liu et al., 2023a), and can be extended to structure-text retrieval tasks and text-based molecule editing tasks.
  • PubChemQA (Luo et al., 2023c). PubChemQA is composed of molecules and their associated textual descriptions sourced from PubChem (Kim et al., 2023). This dataset is characterized by a singular question format: “Please describe the molecule.” This dataset includes 325,754 unique molecules and 365,129 molecule-text pairs, with an average of 17 words per text description.
7.2.2. Protein-Text Datasets
The text-protein dataset focuses on converting text information into protein-related information, including predictions about protein structure, function, and interactions. These datasets typically contain protein sequences, descriptive text, and corresponding annotation information. They have a wide range of applications in bioinformatics and provide a rich resource for the research community to explore new methods and strategies for converting text information into protein information, or vice versa.
  • SwissProtCLAP (Liu et al., 2023f). The SwissProtCLAP dataset, derived from the SwissProt database (Boutet et al., 2007)—a subset of UniProt (Consortium, 2021), consists of 441,000 text-protein sequence pairs. This dataset finds extensive applications in bioinformatics and molecular biology, particularly in tasks requiring the integration of textual descriptions with protein sequences.
  • ProtDescribe (Xu et al., 2023b). The ProtDescribe dataset, comprising 553,052 aligned pairs of protein sequences and corresponding property descriptions, is constructed using data from the Swiss-Prot database (Bairoch and Apweiler, 2000). This database is renowned for its high-quality annotations of diverse protein properties. The dataset facilitates the incorporation of protein property information into protein language models, thereby enhancing their ability to model semantic correlations between known and unknown protein properties.
  • UniProtQA (Luo et al., 2023c). UniProtQA is a dataset comprising proteins and corresponding textual queries about their functions and properties. This dataset, derived from UniProt (Consortium, 2023), encompasses four distinct categories of inquiries related to functional attributes, official nomenclature, protein families, and sub-cellular localizations. It comprises an extensive compilation of 569,516 proteins along with 1,891,506 question-answering instances.
  • InstructProtein (Wang et al., 2023c). The InstructProtein dataset is composed of protein-text instruction data derived from the protein knowledge graph constructed by the UniProt/Swiss-Prot database (Consortium, 2019), which encompasses a comprehensive collection of 2.8 million data. It is the first high-quality protein instruction dataset for tuning LLMs.
7.2.3. Molecule-Protein Datasets
Protein-molecule interaction datasets are an important resource in the field of bioinformatics and drug discovery, providing detailed information on interactions between proteins and small molecule ligands. These datasets are critical for understanding complex interactions between proteins and small molecules, developing new drugs, and predicting molecular behavior.
  • DUD-E (Mysinger et al., 2012). The DUD-E dataset represents an enhanced benchmarking collection developed for evaluating molecular docking. It includes 102 proteins, accompanied by 22,886 clustered ligands sourced from ChEMBL (Gaulton et al., 2012). This dataset was designed to offer a robust and diverse benchmark for well-researched targets, characterized by a substantial number of ligand affinities measurements and a variety of distinct cocrystal ligand structures.
  • BioLiP (Yang et al., 2012). The BioLiP dataset is constructed by utilizing protein structures from the Protein Data Bank (PDB) (Rose et al., 2010) as templates, encompassing a total of 204,223 entries. This dataset serves multiple purposes, including aiding in template-based ligand-protein docking, virtual ligand screening, and annotating protein functions.
  • BindingDB (Gilson et al., 2016). The BindingDB dataset is predominantly constructed using data derived from scientific articles and, to an increasing extent, from US patents. It comprises roughly 1.1 million quantified protein-small molecule affinities, encompassing about 490,000 small molecules and several thousand proteins. This dataset is valuable for various purposes: medicinal chemists can employ it for designing drugs targeted at specific proteins; computational chemists can use it to test and refine algorithms in computer-aided drug design; and pharmacologists can utilize it to screen drug candidates for potential side effects or to formulate hypotheses about the mechanisms of action of new bioactive compounds.
7.2.4. Comprehensive Datasets
The comprehensive datasets amalgamate various data modalities, encompassing protein, molecular, genomic compositions, and associated textual data, thereby enabling a holistic perspective in comprehending biological entities. This dataset plays a crucial role in the training of LLMs for tasks such as prediction of protein-molecule interactions, drug discovery, and automated analysis of biomedical texts, effectively bridging the disciplinary divide between molecular biology and computational linguistics.
  • Galactica (Taylor et al., 2022b). The Galactica dataset, constructed from an extensive scientific corpus, encompasses over 62.22 million documents with 106 billion tokens. It uniquely combines natural language sources, such as academic papers and textbooks, with sequences found in nature, including protein sequences and chemical formulas. Furthermore, this dataset incorporates processed LaTeX to encapsulate intricate scientific data, as well as academic code pertinent to computational science. This composition makes the Galactica dataset an invaluable tool for training neural networks to decode and comprehend the diverse linguistic expressions of scientific disciplines.
  • Scientific Knowledge Dataset (Xie et al., 2023). The Scientific Knowledge Dataset is constructed using two main sources: the SciQ dataset (Welbl et al., 2017), which includes 13,679 science exam questions and answers, and a dataset of 6,057,261 full-text scholarly papers from the Web of Science. This comprehensive dataset is designed to train and enhance large language models for scientific tasks, such as question-answering and knowledge modeling in natural science domains.
  • Mol-Instructions (Fang et al., 2023a). The Mol-Instructions is a comprehensive focus on biomolecular studies and is divided into three main components:
    • Molecule-Oriented Instructions: This segment comprises approximately 1.484 million instructions, distributed across six distinct tasks. It tackles key challenges in various chemical reactions and molecular design.
    • Protein-Oriented Instructions: This segment encompasses roughly 505,000 instructions, spread over five categories of tasks. These tasks are specifically designed to predict protein structure, function, and activity, and also support protein design based on textual instructions.
    • Biomolecular Text Instructions: This portion includes six information extraction and QA tasks represented through 53,000 instructions.
    In summary, the Mol-Instructions dataset is instrumental for the training and improvement of LLMs, specifically in their ability to comprehend and interpret intricate biomolecular data.
7.3. Evaluation
When evaluating MM-Sci-LLMs in the realm of multimodal models, their capabilities are primarily assessed based on three critical areas: cross-modal prediction, retrieval and generation.
Cross-modal Prediction. Cross-modal prediction is a task that involves using multimodal models to predict the functionality of molecules, proteins, and genomes based on textual instructions. These approaches combine different types of data, such as text and molecular structures, to make accurate predictions about the function of biological entities. The specific functions can be found in Section 4.3, 5.3, 6.3. Most encoder-based multimodal models such as MoleculeSTM (Liu et al., 2023a) are designed for this task. Moreover, decode-only architectures such as Mol-Instructions (Fang et al., 2023a) also possess the capability to directly generate functions. As a fundamental task, cross-modal function prediction has various applications in bioinformatics and drug discovery, including identifying potential drug targets, understanding the interactions between proteins, and predicting the activity of molecules.
Cross-modal Retrieval. Cross-modal retrieval refers to the process of retrieving information from one modality based on a query from another modality, and vice versa. In the context of multimodal scientific language models, cross-modal retrieval primarily involves retrieving molecules, proteins, and genomes that are related to textual descriptions, and vice versa. For example, given a textual description of a molecule, a multimodal model can be used to retrieve similar molecules based on their functional descriptions. This can be useful in drug discovery, where researchers may want to find molecules that have specific properties. Models can accomplish these tasks include KV-PLM (Zeng et al., 2022), MoMu (Su et al., 2022) and MolFM (Luo et al., 2023b). On the other hand, given a protein or genomic dataset, a multimodal model such as ProtST-ESM-1b (Xu et al., 2023b) can be used to retrieve textual descriptions that provide information about the properties, functions, or interactions of the proteins or genes in the dataset. This can aid in understanding the biological mechanisms or relationships between different proteins or genes. Additionally, multimodal models can retrieve relevant protein sequences corresponding to a given molecular structure, or find the appropriate molecular structures based on a protein sequence. For example, DrugCLIP (Gao et al., 2023b) uses a dense retrieval approach, employing contrastive learning to maximize similarity between a protein-molecule pair, which enables two-way screening.
Cross-modal Generation. Cross-modal generation tasks aim to generate data in one modality based on data from another modality, which involves establishing interconnections between the different modalities. MM-Sci-LLMs are designed to process data from multiple modalities including texts, molecules, proteins, and genomes. For example, Text2Molecule refers to generating molecular information (e.g., SMILES), based on input text descriptions. Many models can accomplish this task including Text2Mol (Edwards et al., 2021), MolT5 (Edwards et al., 2022), and MOL-Instruction (Fang et al., 2023a). The bidirectional conversion capability of cross-modal generation models allows them to not only generate data in one modality based on data from another modality but also to convert data from one modality to another. For example, Molecule2Text is to produce precise and pertinent textual descriptions derived from molecular data, namely molecule captioning. Similar to the mutual generation of molecule and text, Text2Protein aims to generate protein-related information including sequences and structures based on text descriptions. A representative model is ProteinDT (Liu et al., 2023f) which introduced a text-guided protein design method. Conversely, Protein2Text converts protein sequences into detailed text descriptions to express the complex biological characteristics of proteins. Prot2Text (Abdine et al., 2023), BioMedGPT-10B (Luo et al., 2023c) and MOL-Instruction (Fang et al., 2023a) can address this task effectively. Additionally, Protein2Molecule involves the process of generating relevant molecules from protein sequences, which can be achieved by ChemBERTaLM (Uludoğan et al., 2022) and DrugGPT (Li et al., 2023a). Overall, cross-modal generation in multi-modal models plays a crucial role in bridging the gap between different modalities and enabling the creation of new and cohesive multi-modal data.
Evaluation Metric. The evaluation criteria for MM-Sci-LLMs are similar to those used by uni-modal LLMs introduced in previous sections (Sec. 3.3, 4.3, 5.3, and 6.3), such as the standard classification and retrieval metrics. Hence, we will not elaborate on them again. In particular, in multimodal generation tasks, we emphasize the crucial alignment between the generated content and the input instruction, which necessitates evaluation using specific cross-modal alignment metrics. These metrics evaluate how well the model’s output in one modality (e.g., molecule) aligns with the input or context provided in another modality (e.g., text). However, defining these metrics is challenging due to inherent difficulties in quantifying alignment across different modalities. Therefore, human evaluation often plays a crucial role as expert evaluators can assess biological relevance, novelty, and practical applicability of the generated content, providing insights that automated metrics might overlook.
7.4. Summary
In this section, we conduct a comprehensive examination of the multimodal scientific models MM-Sci-LLMs, which integrate diverse modalities including molecules, proteins, genomes, and textual data. Our investigation entails a detailed analysis of their distinctive architectural designs and capabilities, with particular emphasis on the interactions across multiple modalities. Additionally, we scrutinize the datasets pivotal for training and evaluating these models. MM-Sci-LLMs represent a rapidly growing research direction that integrates multiple modalities of scientific data. Currently, this field is still in its early developmental stages, offering significant opportunities for pioneering works and innovative applications in science discovery.
8. Conclusion and Perspective
In this survey, we have explored the latest advancements in Sci-LLMs, particularly in the biological and chemical fields. The survey began by establishing the foundational concepts of scientific languages, encompassing textual, molecular, protein, and genomic categories, as detailed in Sec. 2. We then delved into each scientific language type, examining their recent developments. This examination included an in-depth look at their model architectures, capabilities, datasets, and evaluation, covered in Sec. 3, Sec. 4, Sec. 5, and Sec. 6, respectively. A notable highlight of the survey is Sec. 7, where we presented the emerging multimodal Sci-LLMs, focusing on the integration of multiple scientific language systems.
In what follows, we discuss the research challenges of Sci-LLMs and potential avenues for future exploration. Given the considerable advantages and growing prominence of generative LLMs in scientific applications, our discussion will primarily focus on these generative Sci-LLMs. This targeted approach allows us to thoroughly examine the unique aspects and possibilities that generative models offer in advancing the field of Sci-LLMs.
8.1. Critical Challenges
Although previous research has made notable strides in the field of Sci-LLMs, it is important to acknowledge that this area of study is still in its nascent stages. In the process of preparing this survey, we conducted a thorough examination of existing studies and identified several critical challenges that are yet to be addressed. These challenges highlight the evolving nature of Sci-LLM research and underscore the need for continued exploration in this domain.
8.1.1. Training Data
Data stands as the foundation of AI model development. In exploring Sci-LLMs, we focus on key factors that influence their growth and effectiveness.
  • Scale of Pre-training Datasets. The scale of pre-training datasets is a pivotal factor constraining the advancement of larger-scale Sci-LLMs. In the realm of language models, scaling laws delineate the relationship between an LLM’s size (quantified by parameters or computational resources) and its performance, indicating that a model’s capabilities and effectiveness tend to enhance, often in a nonlinear or exponential manner, as its size increases. Given that data acts as the fuel for AI models, amassing a substantial volume of pre-training data is critical for the development of larger and potentially more proficient LLMs. When contrasting general LLMs with Sci-LLMs, a stark difference emerges in their data scales. General LLMs often comprise hundreds of billions of parameters, while Sci-LLMs typically operate on the scale of billions. This disparity is largely due to the vast abundance of human language data on the web. For instance, ChatGPT, with its 175 billion parameters, is estimated to be trained on 570 billion tokens. In contrast, a protein language model like ProGen, despite having 1.2 billion parameters, is limited to training with 280 million protein sequences, encompassing tens of billions of tokens. However, there is a silver lining. With the continuous advancements in sequencing technologies, including DNA and protein sequencing, more gene and protein sequences are being identified. These advancements are expected to significantly contribute to the development of larger and more capable Sci-LLMs.
  • Quality of Finetuning Datasets. Following pre-training, Sci-LLMs undergo a supervised fine-tuning process using labeled datasets. This stage is pivotal as it trains the models to follow specific instructions and generate the desired outputs. The effectiveness of this fine-tuning is heavily dependent on the quality of the labeled datasets, which can be assessed in terms of scale, diversity, and balance. A larger scale of fine-tuning datasets generally leads to more proficient models, as they are exposed to a wider array of scenarios. However, compiling extensive annotated data can be resource-intensive. Besides, data diversity is also crucial; reports have shown that exposing LLMs to a broader spectrum of instructions enhances their ability to generalize across various downstream tasks. Equally important is the balance within the datasets. An imbalanced distribution of labels can lead Sci-LLMs to disproportionately favor predictions in the majority class, undermining their effectiveness.
  • Lack of Cross-modal Datasets. The quality of these cross-modal datasets, predominantly used in the fine-tuning stages, must be rigorously assessed in terms of scale, diversity, and balance. However, beyond these factors, the alignment between modalities emerges as a crucial and more complex consideration. Alignment entails ensuring that each data point contains samples from different language systems which express the same or relevant semantic content. This requirement demands consistent semantic representation across different modalities. This deficiency could result in a narrowed scope of instructional diversity and a limited span of language systems included. To foster the effective development of robust multi-modal Sci-LLMs, it is essential to ensure that these datasets are not just extensive in scale but also rich in diversity and balance. They must also maintain strong semantic alignment across the various modalities, thereby enabling more accurate cross-modal interactions and interpretations.
The extensive pre-training corpus of Sci-LLMs provides vital exposure to a diverse array of semantic and linguistic elements. In fine-tuning these models, it is imperative to focus on aspects such as scale, diversity, balance and alignment across modalities. Such attention ensures the maximization of Sci-LLMs’ performance in various applications.
8.1.2. Architectures and Learning Objectives
While most LLMs rely on Transformer-based architectures to learn semantic relevance in languages, this approach may not be ideally suited for Sci-LLMs. Several reasons underpin this mismatch:
  • Handling Longer Sequences. Scientific languages often comprise sequences that are significantly longer than typical natural language sentences. For instance, a protein sequence might contain hundreds or thousands of amino acids, while a natural language sentence usually has fewer than 30 words. This disparity in sequence length poses computational efficiency challenges for traditional Transformer-based architectures. To address this, recent models such as RWKV, Hyena, and S4 have been developed, showcasing innovations in handling longer sequences more effectively
  • Incorporating 3D Structural Information. Current biological and chemical language models struggle to explicitly incorporate crucial 3D structural information, which is often more important than sequences alone for understanding functions and properties. 3D structural data, typically represented as atom coordinates, does not seamlessly integrate into traditional language systems. Recent studies have attempted to map 3D structures to learning tokens, aiming to create structural tokens in addition to sequence tokens. These efforts, though promising, are still in their infancy.
  • Autoregressive Learning Objective Limitations. Natural languages are generally processed in an autoregressive manner. However, this is not the case with biological and chemical languages, which are not inherently created or interpreted autoregressively. In these domains, sequences function as a whole. Therefore, an ideal learning objective for a Sci-LLM should allow the model to capture semantic information from the entire sequence, considering bidirectional context, rather than relying solely on one-way information flow.
These considerations highlight the need for specialized adaptations or entirely new model architectures and learning objectives for Sci-LLMs to effectively process and understand scientific languages, taking into account their unique characteristics and requirements.
8.1.3. Model Evaluation
The effective evaluation of LLMs remains a critical area of research focus. As detailed in Sec. 3, researchers often utilize Bloom’s taxonomy, a framework defining six cognitive levels of educational objectives, to assess learning and understanding in LLMs. However, the unique nature of Sci-LLMs necessitates a tailored approach to evaluation. Unlike general LLMs, Sci-LLMs are primarily geared towards mastering scientific knowledge and demonstrating utility in research-oriented scenarios. To address this, we propose a novel evaluation framework that categorizes Sci-LLM capabilities based on the complexity of scientific knowledge and research acumen. This framework classifies capabilities into pre-college, college, and post-college levels, aligning the model’s understanding and output with the sophistication expected at these educational stages.
However, this methodology predominantly addresses textual outputs. The challenge intensifies when evaluating the generation of biological and chemical languages, such as those used in the de novo design of proteins or molecules. Currently, it is arduous for computational systems, and even more so for human experts like biologists and chemists, to accurately ascertain the characteristics and efficacy of these generated entities based solely on observation. Computational methods like molecular simulation offer some insights, but their reliability can vary, sometimes deviating significantly from actual outcomes, not to mention that they are usually too slow to closely collaborate with Sci-LLMs.
The computational metrics discussed in Sec.4, Sec.5, and Sec. 6 offer indirect indications of the quality of newly generated proteins or molecules. However, these indicators are not always definitive. Wet-lab experiments remain the gold standard for validation, yet they are often beyond the scope and capabilities of many AI research teams. This gap underscores the urgent need for more robust and reliable computational evaluation and benchmarking systems. Such systems would not only augment the accuracy of assessments in the absence of wet-lab experimentation but also enhance the overall development and application of Sci-LLMs in scientific research.
8.1.4. Ethics
The ethical considerations in the development and application of Sci-LLMs are multifaceted and critical. Firstly, data privacy and consent are paramount, especially when dealing with sensitive biological data. Ensuring that genomic data, for instance, is anonymized and used with proper consent is vital to maintaining individual trust. Secondly, there is the risk of misuse of information, such as the potential for creating harmful biological substances or contributing to bioterrorism, necessitating stringent control measures. Bias in algorithmic decision-making also poses a significant ethical challenge. Sci-LLMs, like general LLM systems, can perpetuate and amplify biases present in their training data, leading to skewed outcomes in scientific research and applications. Lastly, there is an ethical imperative to ensure equitable access to the benefits of Sci-LLMs, preventing the exacerbation of existing inequalities in scientific research and healthcare. Addressing these ethical issues requires a collaborative approach involving ethicists, scientists, policymakers, and other stakeholders to develop robust guidelines that ensure responsible and beneficial use of Sci-LLMs.
8.2. Future Directions
To propel forward the field of AI-driven scientific discovery, we have pinpointed five key research avenues that hold promise for the advancement of more capable Sci-LLMs in the future.
  • Construction of Larger-scale, High-quality, and Cross-modal Training Datasets. As we analyzed before, a critical challenge lies in the realm of training data. Expanding the pre-training dataset is contingent upon advancements in biological sequencing technologies and the open-sourcing of chemical molecular data. Concurrently, the responsibility of maintaining high-quality fine-tuning datasets, particularly in the cross-modal domain, falls to AI research teams. These enhanced datasets are instrumental in not only improving the understanding of single-modal language systems by Sci-LLMs but also in enabling them to tackle biological and chemical tasks from a more integrated, cross-modal perspective. This dual focus on expanding and refining training datasets is crucial for realizing the potential of larger and more effective Sci-LLMs in practical scientific research scenarios.
  • Incorporating 3D Stereo-Temporal Information into Scientific Language Systems. The inherent nature of biological and chemical data is rooted in the three-dimensional physical world, where the functionalities are primarily determined by their constantly changing 3D structures. This highlights the urgent need to integrate 3D structural and temporal information into language-based modeling approaches. In addition to the recent structural tokens, it may be equally crucial to explore structural motifs. These motifs, akin to Lego building blocks in their ability to interlock and compose protein conformations, represent specific 3D structures that hold significant potential for modeling. By incorporating these structural motifs into the scientific lexicon, they can be learned and processed by Sci-LLMs in a manner similar to other tokens. Besides, the learning process is supposed to capture the dynamic alternation of 3D structures. This would enrich the models’ understanding and representation of the intricate three-dimensional aspects of biological and chemical data.
  • Sci-LLM’s Marriage with External Knowledge Sources. After centuries of development, scientists have accumulated a wealth of domain knowledge. For instance, biologists have developed the Gene Ontology, a comprehensive knowledge graph that systematically delineates the relationships between protein sequences and their functions. The potential synergy between Sci-LLMs and these external knowledge sources represents a substantial opportunity for enhancement. By integrating Sci-LLMs with such rich, structured external resources, their capabilities can be significantly expanded, overcoming the potential issues such as hallucination. This approach underscores the importance of looking beyond sequence data and embracing a more holistic view in the development of Sci-LLMs.
  • Sci-LLM Interacting with Physical Simulation. Current Sci-LLMs are primarily adept at capturing the semantics inherent in sequence data. However, the realm of scientific understanding extends far beyond sequence data alone, encompassing a wealth of additional resources that can significantly contribute to a deeper comprehension. For instance, the field of Molecular Dynamics (MD) offers simulations that mirror the physical world, providing insights into molecular behavior and interactions. The interplay between Sci-LLMs and MD would not only complement the existing data-processing strengths of Sci-LLMs but also enrich their analytical power, leading to more nuanced and accurate interpretations of complex scientific phenomena.
  • Augmenting Sci-LLMs with Specialized Tools and Agents. General LLMs have demonstrated remarkable proficiency in leveraging software and tools, such as calculators, to accomplish complex tasks. This capability is equally pertinent to Sci-LLMs, which can be significantly enhanced through integration with science-specific computational tools. A prime example of this is the recent success of GPT-4, which has shown proficiency in web searching, academic paper analysis, and designing experiments, including replicating intricate cross-coupling reactions like the Suzuki and Sonogashira experiments. The realm of scientific tools is vast and intricate, encompassing a wide array of computational resources. The challenge of accurately selecting the appropriate tool or software from tens of thousands of possibilities, precisely inputting data, and effectively interpreting the output, presents a rich area for research exploration. The advancement in this area holds significant promise for augmenting the capabilities of Sci-LLMs, enabling them to perform more sophisticated and specialized scientific tasks with higher precision and efficiency.
  • Development of Computational Evaluation Metrics and Benchmarks. The proper evaluation of model performance is crucial, especially for generative tasks. This extends beyond the capabilities of AI teams alone, necessitating greater collaboration with experts in computational biology. Unlike the time-intensive process of molecular dynamics, the ideal evaluation metrics for these models should be both rapid to compute and accurate in reflecting outcomes. Establishing such metrics would facilitate the quick construction of evaluation benchmarks, significantly speeding up research iterations. This collaborative approach, merging AI with computational biology, is essential for developing effective, efficient evaluation strategies that can keep pace with the rapid advancements in model development.
  • Super-alignment with Human Ethics. Alignment with human ethics is pivotal in the development of Sci-LLMs. Super-alignment, in this context, extends beyond basic adherence to ethical norms and encompasses a deep integration of ethical reasoning capabilities within Sci-LLMs. This involves equipping these models with a nuanced understanding of ethical principles and the ability to apply them contextually in diverse scientific domains. For instance, in genomic research, Gene-LLMs should be able to weigh the benefits of genetic discoveries against potential risks like genetic discrimination or privacy breaches. Future research in this area could focus on developing frameworks and algorithms that enable Sci-LLMs to recognize and respect diverse ethical viewpoints, allowing them to suggest or evaluate actions based on ethical principles. Ultimately, the goal is to ensure that the advancements driven by Sci-LLMs are both scientifically robust and ethically sound, thereby fostering trust and acceptance in their applications across various scientific fields.
In this survey, we have conducted a systematic review of scientific large language models, particularly emphasizing the biological and chemical domains. We established the foundational concepts of scientific languages including scientific texts, molecules, proteins, and genomes. Then we delved into each scientific language type, examining the recent developments in Sci-LLMs. This examination included an in-depth look at their model architectures, capabilities, datasets, and evaluation methods. Moving forward, we pinpointed four key challenges and proposed seven promising research directions for future exploration. Our goal is to serve as a comprehensive and insightful resource for both the AI and fundamental science communities, fostering collaborative efforts and driving forward the ‘AI for Science’ research agenda. By effectively modeling scientific languages, LLMs pave a more stable path towards the realization of artificial general intelligence.
Contributions
This survey is led and organized by Qiang Zhang, Keyan Ding and Huajun Chen, and the contributions of all authors for each section are listed as follows:
  • Sec.1: Qiang Zhang, Keyan Ding, Huajun Chen
  • Sec.2: Qiang Zhang, Keyan Ding, Renjun Xu, Huajun Chen
  • Sec.3: Jing Yu, Keyan Ding, Qiang Zhang, Zhuoyi Xiang, Huajun Chen
  • Sec.4: Tianwen Lyu, Keyan Ding, Qiang Zhang, Xiang Zhuang, Jiyu Cui, Huabin Xing, Huajun Chen
  • Sec.5: Yiwen Zhang, Keyan Ding, Qiang Zhang, Zeyuan Wang, Hongyang Chen, Huajun Chen
  • Sec.6: Qingyu Yin, Keyan Ding, Qiang Zhang, Mengyao Zhang, Jinlu Zhang, Xiaohui Fan, Huajun Chen
  • Sec.7: Xinda Wang, Keyan Ding, Qiang Zhang, Xiaotong Li, Yuhao Wang, Ming Qin, Huajun Chen
  • Sec.8: Qiang Zhang, Keyan Ding, Huajun Chen